### v1

# Evaluating link prediction: New perspectives and recommendations

Abstract
Link prediction (LP) is an important problem in network science and machine learning research. The state-of-the-art LP methods are usually evaluated in a uniform setup, ignoring several factors associated with the data and application specific needs. We identify a number of such factors, such as, network-type, problem-type, geodesic distance between the end nodes and its distribution over the classes, nature and applicability of LP methods, class imbalance and its impact on early retrieval, evaluation metric, etc., and present an experimental setup which allows us to evaluate LP methods in a rigorous and controlled manner. We perform extensive experiments with a variety of LP methods over real network datasets in this controlled setup, and gather valuable insights on the interactions of these factors with the performance of LP through an array of carefully designed hypotheses. Following the insights, we provide recommendations to be followed as best practice for evaluating LP methods.
keywords:
Link prediction, Evaluation, Experimental setup,Early retrieval, Class imbalance1 Introduction
Complex networks encode rich and non-trivial relations among entities. They are ubiquitous in nature, and are used to model many complex real-world systems. Social, biological and information systems are to name a few such systems. Social networks encode some kind of social relationships, biological networks model interactions between two biological entities like proteins, and information networks record information exchanges, such as emails. Each entity in a complex network is called a node and any connection, interaction, or exchange of information between two nodes are referred as a link. Graph data structure is used to encode complex networks, where vertices represent nodes and edges represent links. Link prediction (LP) [1, 2, 3] is a fundamental problem in network science, which, given a network, aims at predicting unobserved, missing, hidden or future relationships, depending on the complex system and the target application [4]. LP has numerous applications: it can suggest unobserved or missing relations between two criminals in a criminal network [5, 6], it can predict future collaborations among scientists from co-authorship networks [7, 8, 9], it can recommend unobserved costumer-product relationships in collaborative filtering [10, 11, 12], and so on. LP has been primarily targeted as a classification problem [2] which, given a network, classify a given node pair, which are not connected by a link in the given network, to yes or no: yes signifying they have a hidden connection or they will be connected in the future, and no means there is no possibility of getting connected by a link.
Since Nowell et al. [1] popularized LP as a research problem, now it is a well established research domain in both of network science and the machine learning community. Application of LP is diverse, so a solution to this problem demands considerations of the application or the domain specific factors. Although, there exists application specific designs of the LP methods [13, 14, 15, 11, 16, 7], they usually extend the state-of-the-art node similarity [1, 2, 17] or machine learning based [18, 19, 20, 21, 22, 23, 11] methods. These state-of-the-art methods assume a uniform set-up accross different prediction cenerios, starting from problem formulation to experimental setup and evaluation. We investigate if this uniformity hinders accurate evaluation of the LP methods, and aim to provide a guideline for a robust evaluation of LP. We identify a few aspects concerning the data, experimental setup and evaluation metric that may affect the prediction performance, based on which we build our experimental framework, and perform extensive experiments to gain insights toward effective and robust evaluation for LP. We brief those aspects below.
The LP problem can be categorized into two: (a) missing/unobserved/hidden link prediction, and (b) future link prediction [1, 2, 24]. Without loss of generality, now onward we refer these two types as missing LP and future LP respectively, and name this classification as prediction type. Traditionally, in literature, the experimental setups for LP do not differentiate between these two prediction types, and evaluate LP in a common experimental setup which ignores the information of the link appearance time, and the ground truth node pairs are generated by removing a fraction of random edges from the underlying graph. This results in information loss and makes the future LP problem get evaluated as a missing LP problem. This may not capture true efficacy of a prediction method on future LP, because a randomly deleted link may not be a true future link. Our experimental setup accommodates both types of LP, and we examine if this information loss results in any performance degradation of the LP methods. A line of works target temporal link prediction [7, 25], which devices specialized LP methods by exploiting temporal dynamics in dynamic networks. In this work, we are not focusing on temporal link prediction, but consider the traditional approach of solving LP.
Popular LP methods are of two types: similarity based and machine learning based [24]. Similarity based methods [1] calculate some score based on structural similarity or closeness of a node pair, which measures the likelihood of having a link. These methods can be either local or global. Local methods rely on triadic closure property [26] which says that two unconnected nodes tend to form a link if they have at least one common connection. Local methods are applicable on a pair of nodes whose shortest path distance in the network graph, which is also referred as geodesic distance in network science, is two. To increase readability, we refer “geodesic distance” as “distance” throughout this paper. We refer a node pair having distance two as two-hop away/apart node pairs. Global methods are mostly shortest path or random walk based, which can predict connections between any node pair in the network. Modern machine learning based LP methods either train a neural network in a supervised way [18, 27, 28, 29], or generate low dimension embedding for the nodes in unsupervised manner and use the embeddings for predicting links [19, 20, 21, 22]. Machine learning based methods are global in nature. The small world property [30] makes the diameter of a complex network grow at a logarithmic rate with network size, and the likelihood of connection decreases with the distance between two nodes. The LP methods which are global in nature, mostly agree with this principal. So, in one hand, the local methods apply on two-hop away node pairs, whereas the global methods apply on two-hop as well as more distant node pairs; on the other hand, the new connection build-up process is biased toward the two-hop away node pairs than the rest. Traditional evaluation frameworks do not distinguish the test node pairs by the distance of the end nodes in the network. So, the evaluation does not capture the unique characteristics, if any, of the test node pairs based on the distance of the end nodes. Moreover, when we compare a global method with a local one in this framework, we may draw wrong conclusions because the test set for the two cases may differ. Also, we miss the opportunity to compare two global methods in a distance controlled setup. The traditional evaluation approach has another inherent issue. The negative test examples are sampled by selecting random node pairs which are not connected by a link in the network, and are not in the positive examples. This randomness makes the distribution of negative examples differ with the positive ones, for which the samples have an inherent distance bias. We propose an experimental set-up which takes care of these factors, and conduct an array of systematic experiments to find valuable insights toward effective evaluation.
The state-of-the-art LP methods usually apply on undirected networks. However, networks like information networks are directed in nature. So, these networks are usually converted to undirected networks by ignoring their directions in order to make those methods applicable on them. Ignoring directions results in loss of significant information about the links, which may affect the prediction performance in directed networks. We investigate this with experiments in our hop controlled and prediction type controlled setup.
Class imbalance is an inherent issue with LP, where the real missing or future links are too few as compared to the node pairs with no possible connections. Historically, Area Under the Receiver Operating Characteristics Curve (AUROC) [31] has been a popular single point summary statistics evaluation metric for LP [2, 24], perhaps due to its invariance with class distribution. AUROC favors accurate classification of positive links, at the cost of significantly misclassifying negative node pairs. Hence, AUROC can overestimate the performance of LP. Skew sensitive measures like Area Under the Precision-Recall Curve (AUPR) or Precision@k can be an alternative [32, 3, 2], but use of these are rare in literature, may be due to the difficulty in handling the variability in results with skew level for their skew sensitiveness. So, in order to understand how an LP method behaves in skewed test environment, we require a rigorous skew controlled evaluation with appropriate evaluation metrics. Moreover, applications like recommendation systems require methods to perform good at the early retrieval stage. Early retrieval performance is skew sensitive, the effect of which can not be understood with AUROC, or the ROC curve [33, 34, 32]. So, how the LP methods behave in the early retrieval stage in imbalanced setup has been largely unexplored. Our proposed experiment setup and the conducted experiments take care of these factors and judiciously employ the evaluation metrics to understand the true behavior of the prediction methods considering class imbalance and early retrieval. It also examines the interplay of the other factors like prediction type and distance between the end nodes of the test node pairs with class imbalance.
Our contributions in this paper can be summarized as follows.
we identify a number of aspects concerning the data, experimental setup and evaluation metric that may affect the performance of LP methods, which the baseline setups for the LP evaluation largely ignore;
we propose an experimental setup which allows us to evaluate LP methods in a rigorous manner in a controlled environment, considering those aspects;
we conduct extensive experiments using 24 real network datasets on 58 LP methods in this controlled experimental setup, and perform an array of correlation and hypothesis tests to analyze the effect of those factors on LP methods;
we provide recommendations to be followed as best practice toward an effective and robust evaluation of LP methods.
This paper is organized as follows. In Section 2, we summarize the related works. We provide the details of the LP methods which we use in our experiments in Section 3, grouped by their categories. Section 4 provides the details of the datasets on which we apply the LP methods. We describe the proposed experimental setup in Section 5. In Section 6, we present the results and analyze our findings, and provide our recommendations. We conclude our work and indicate the future directions in Section 7.
2 Related Works
Nowell et al. [1] proposed a list of local and global similarity based LP methods. Similarity based methods calculate a score between two nodes to estimate their chance to form a link. Local similarity based methods are neighborhood based, and apply to two-hop away node pairs. The global similarity based methods are mostly path based and apply to any node pair in the network. In this paper, we consider most of the methods proposed in Nowell et al. in our experiments along with another popular local method: Resource Allocation [17]. There are a few other similarity based methods in literature, an exhaustive list of which can be found in Kumar et al. [2]. We do not include another type of similarity based method, namely quasi-local [2], in our study. Modern machine learning based LP methods are either supervised [18, 27, 28, 29] or unsupervised [19, 20, 21, 22]. The supervised methods train a neural network or a graph neural network in an end-to-end manner with the test node pairs. The unsupervised methods first train a neural network or a graph neural network to learn low dimension embeddings of the nodes in an unsupervised manner (known as contrastive learning), and then produce link features which are used to build a classifier to predict links. In this paper, we consider the unsupervised methods in our experiments. The main reason to exclude the end-to-end methods is the difficulty in building the experimental set-up for those in the longitudinal setup for the future LP, which incurs a longitudinal bias [3, 35] on the prediction performance, as well as makes the methods incomparable with their missing LP counterpart. Another group of unsupervised LP approaches are matrix factorization based [23, 36], which we do not include in our experiments. Note that, the purpose of this work is not to present an exhaustive performance comparison of the existing methods, but to devise a robust evaluation strategy for the prediction methods considering certain data, evaluation metric, experimental setup, application and methodology specific factors. So, excluding a few prediction methods does not hurt the merit of this work.
The start-of-the-art LP methods briefed above mainly apply to a simple, undirected and homogeneous networks. There exists works which exploit several network attributes to predict links. Few examples are: temporal link prediction [7, 25] which exploits temporal dynamics of link maintenance in dynamic networks like in social networks; link prediction in bipartite networks [37, 38] which applies on bipartite graphs like user-product networks or term-document networks; link prediction in heterogeneous networks [7, 39, 40] which exploits the heterogeneity of nodes and links in networks like bibliographic networks; link prediction in directed networks [41, 42] which exploits the direction information in directed networks like email networks, etc. These specialized methods are not necessarily disjoint, and often use information covering multiple groups. In this paper, we consider the start-of-the-art LP methods only in our analysis, and reserve the specialized methods for future study.
Studies like [43, 35] investigated the issues associated with LP evaluation in considerable details. They stressed on an imbalanced sampling of test set, and advocated for AUPR and the PR curve to tackle class imbalance in evaluation. Yang et al. [35] also proposed a test set-up for supervised future LP from deployment perspective, and analyzed the effect of variation of the test node pair collection time duration. He et al. [44] investigated how LP accuracy in real-world networks gets affected under non-uniform missing-edge patterns. A few studies like [45, 46] investigated the evaluation problem for temporal link prediction. These studies consider the timestamps of test node pairs as an important aspect for evaluation. Kumar et al. [2] presents a nice survey on LP methods and evaluate them using standard metrics like AUROC, AUPR and Precision@K. Mostly, the existing studies on link prediction evaluation deal with the class imbalance issue [43, 35, 2, 3, 47, 48], but do not address the early retrieval performance. To the best of our knowledge, there exists no study which considers the multifaceted aspects of LP evaluation in such a detailed manner like us.
3 Link Prediction Methods
3.1 Similarity Based Methods:
Similarity-based LP methods calculate a score given a pair of nodes, which represents the likelihood of a link formation between them. Based on the node pair’s geodesic distance in the network, similarity based methods are classified into two: local and global. Local methods are applied on the node pairs which are two-hop apart, but global methods can be applied on any node pair irrespective of their distance in the network. As the machine learning based methods also are global in nature, for ease of understanding, we refer local similarity based methods as local-sim and global similarity based methods as global-sim now onward. We summarize a few popular local and global similarity based methods which we use in this paper as follows.
3.1.1 Local methods
Common Neighbors (CN) [1]: The Common Neighbors method (CN) measures the number of common neighbors or two-hop paths between node pairs. The score between nodes and can be expressed as:
where represents the set of neighbors of . Jaccord’s Coefficient (JC) [49, 1]: Jaccord’s Coefficient (JC) extends the method by penalizing it for non-shared neighbors between the nodes. The score between nodes and can be expressed as:
Adamic Adar (AA) [1, 50]: Adamic Adar (AA) extends the method by penalizing each common neighbor by its degree logarithmically. The score between nodes and can be expressed as:
Resource Allocation (RA) [17]: Unlike , Resource Allocation (RA) penalizes each common neighbor by its degree without logarithmic scaling. The score between nodes and can be expressed as:
3.1.2 Global methods
Preferential Attachment (PA) [51, 52, 1]: Preferential Attachment method (PA) relies on the principle that nodes with higher degrees are more likely to acquire new connections. The score between nodes and can be expressed as:
Katz Similarity (Katz) [53, 1]: Katz similarity index (Katz) enumerates all the possible paths of different lengths between the node pairs, and takes sum over this collection, exponentially damping by path length. The Katz score between nodes and can be expressed as:
where acts as decay factor to give exponentially higher weight to longer paths, and is the number of different paths of length connecting the node pair. Hitting Time (HT) and Normalized HT (Norm-HT) [1]: Hitting Time (HT) leverages random walks on a network to quantify node similarity. The hitting time between nodes and is the expected number of steps it takes for a random walker starting at node to reach node for the first time. It quantifies how easily information or influence can spread between the nodes. It indicates easier information flow or shorter travel times between the nodes in the network, where smaller HT indicates better link formation likelihood. The scoring function can be expressed as:
Here, is the random variable denoting the time it takes for a random walker to reach node from , and is the probability of this happening in steps. is quite small when the node has a large stationary probability. To mitigate this issue, the score is multiplied with ’s stationary probability. We call this measure as normalized heating time (Norm-HT).Commute Time (CT) and Normalized CT (Norm-CT) [1]: Commute Time (CT) signifies the expected time for a random walker to travel from node to and back to . It encapsulates the symmetric nature of node connectivity. The CT score between nodes and can be expressed as:
Like HT, we consider normalized commute time (Norm-CT) along with its un-normalized version.
3.2 Machine Learning Based Methods
We use three popular node embedding methods in this category, namely, Deepwalk [19], Node2vec [20], and Graphsage [21], towards LP. All of these methods learn a function , given a network graph , where and are the set of vertices and edges respectively, which maps each node to a dimensional latent space where . We learn edge features following the method presented in Grover et al. [20] to get a link vector given a node pair, and apply logistic regression and random forest supervised learning technique to predict links. We refer this group of LP methods as learning. Below, we brief the aforementioned node embedding methods and the edge feature learning methods.
3.2.1 Deepwalk [19]:
Deepwalk adapts Skip-gram [54] method of natural language processing to generate node embeddings. It solves the following optimization problem:
where are the context vertices of the vertex within a window of size . To get the context vertices, Deepwalk simulates multiple truncated random walks started at every vertex, and extracts the context vertices from the walks. It approximates the optimization by solving it with a conditional independence assumption on the vertices in the context window, given the embedding of a vertex. On this assumption, it models the conditional probability of each node-context pair with a softmax unit. The optimization equation can be given by:
where , and are the set of nodes inside the context window of . To avoid the explosion of labels, precisely numbers, Deepwalk uses Hierarchical Softmax [55, 56] with stochastic gradient descent (SGD) to approximate the optimization.
3.2.2 Nodevec [20]:
Like Deepwalk, Nodevec adapts Skip-gram to generate the embedding, and solves a similar optimization problem. It differs from Deepwalk in two aspects: generating the context window, and approximating the optimization problem. It argues that embeddings generated using truncated random walk can not capture homophily and structural equivalence among nodes. In order to capture a more flexible contextual structure, it employs a biased random walk method that combines neighborhood exploration using breadth-first search (BFS) and depth-first search (DFS). The context sampled by BFS generates similar embeddings for structurally equivalent nodes. DFS can explore distant parts of the network, which leads the embeddings to preserve homophily. Like DeepWalk, with the conditional independence assumption over the context window, Nodevec models the conditional probability with a softmax units. Nodevec deals with the explosion of labels by negative sampling with SGD to approximate the optimization.
3.2.3 GraphSAGE [21]:
GraphSAGE is a graph neural network approach for scalable and inductive learning on large graphs. GraphSAGE leverages node attributes (e.g., node2vec embeddings) to learn embedding functions that generalize to unseen nodes during the training phase. GraphSAGE does this by learning aggregator functions that can induce an embedding of a node by aggregating the attributes of its neighboring nodes, sampled from its direct connections. This aggregation process is executed times for all nodes in the network, which way it learns sets of weight matrices . It uses four different aggregator techniques: mean, max-pooling, mean-pooling, and LSTM. GraphSAGE optimizes similar objective as Deepwalk and Nodevec, and approximates it with negative sampling, where given a node, its positive instances are sampled from nodes appearing in the chain of short random walks starting at the given node, and negatives are sampled from the degree distribution. It uses SGD as the optimization procedure.
3.2.4 Learning edge features:
To learn edge features, we follow the procedure described in Grover et al. [20]. For two nodes and , we employ a binary operator “” for the feature vectors and in order to achieve an edge representation such that where is the representation size for the pair . The operators designed for this purpose are designed to be adaptable, allowing them to be applied to any node pair, even if there is no existing edge. This flexibility makes representations more valuable, especially in LP settings when there are both true and false edges in the test set. We summarize the operators in Table 1.
| Operator | Symbol | Definition |
| Average | ||
| Hadamard | ||
| Weighted-L1 | ||
| Weighted-L2 |
3.2.5 Naming convention of individual link prediction methods:
Now onward, we refer each of the similarity based methods with their abbreviations. There are multiple variants of GraphSAGE for different aggregation techniques. There are multiple variants of the LP methods using each of the embedding methods depending on the edge learning operator, and the supervised learning method used. We abbreviate individual LP methods generated from these embedding methods as: (embedding-method)_(aggregation-technique-if-applicable)_(edge-learning-operator)_(supervised-method). To improve presentability, we further abbreviate Deepwalk, Nodevec and GraphSAGE as DW, NV and GS respectively, the aggregator techniques mean, max-pooling, mean-pooling and LSTM as Mean, MeanPool, MaxPool and LSTM respectively, the edge learning operators Average, Hadamard, Weighted-L1 and Weighted-L2 as Avg, Hada, L1 and L2 respectively, and the supervised learning methods logistic regression and random forest as LR and RF respectively. For an example, GS_Mean_Avg_RF refers to the LP method which generates the node embeddings using Mean aggregator technique with GraphSAGE, and then learns edge features using the Average operator, and then trains a random forest model to predict links. In total, our experiments are conducted on 58 distinct LP methods, of which, 4 are local-sim, 6 are global-sim, and 48 are learning based methods.
4 Datasets
We conducted our experiments on fifteen real network datasets of diverse nature, such as, social network, communication network, collaboration network, biological network, power grid, etc. A few of these networks like friendship networks, collaboration networks are undirected in nature, whereas networks like communication networks are directed. The social, communication and collaboration networks are inherently dynamic or temporal in nature. Among the networks in our collection which are dynamic, few had timestamps available with the interactions or the moment of link formation. In total, nine among all the networks had timestamps. For these networks we created ground truth labels two ways: (a) considering the timestamps which lets us evaluate future LP, and (b) removing a portion of the links in the networks which lets us evaluate the missing LP. So, we created two LP datasets for each of these networks. For the networks with no timestamp, we could only evaluate missing LP. In total, we conducted our experiments and analysis on datasets, among which 15 were used for missing LP and the rest were used for future LP. We summarize the datasets in Table 2. We briefly describe the datasets as follows.
4.1 Networks with timestamps:
1.
CollegeMsg [57]: This dataset represents a temporal network capturing private communications among college students on a campus at the University of California, Irvine. Each node in the network represents an individual student, while the private messages exchanged between them guides the link formation and maintenance. This dataset offers insights into communication dynamics and prevalent styles within the campus community. Information for this dataset was obtained over the course of 193 days at the University of California, Irvine. In this paper, we refer the dataset which prepares ground truth using timestamps as clg-msg-t and the the one which do not consider the timestamps as clg-msg.
2.
Email-Eu [58]: This dataset provides insight into the evolving connections and communication patterns within a European research organization over time. It primarily focuses on email correspondence exchanged between members of the institution. The network’s nodes represent individual users or email addresses within the organization, and the links encode the email exchanges between them, with timestamps indicating the exact moments of interaction. This data enables researchers to analyze the dynamics of information flow and collaboration within the organization. Information for this dataset was gathered over a period of 803 days. In this paper, we refer the temporal version as email-t and the non-temporal one as email.
3.
Mathoverflow [58]: This dataset presents a temporal network capturing interactions among users on the Mathoverflow platform, a specialized community for mathematicians to exchange knowledge and insights. Each node in the network represents an individual user, while the links encode their interactions, such as comments or responses to questions and answers. The timestamps associated with the interactions indicate the exact moment of these interactions, providing valuable temporal insights into the dynamics of knowledge sharing and collaboration within the mathematical research community. This dataset is taken over a time span of 2350 days. In this paper, we refer the temporal version as math-flow-t and the non-temporal version as math-flow.
4.
Ast-phys [59]: This dataset contains a collaboration network of authors who have contributed to the field of Astrophysics through scientific paper publications. Each node represents an individual author, and the links signify collaborations between authors. The timestamps indicate the time of collaborations between two authors. In this paper, we refer the temporal version as ast-phys-t and the non-temporal version as ast-phys.
5.
LastFm [60]: This dataset represents a social network derived from Last.fm, a music streaming and recommendation service. Nodes in the network represent Last.fm users, and the links symbolize their interactions and social relationships. The dataset includes timestamps indicating the precise moments of user connections and interactions. In this paper, we refer the temporal version as lst-fm-t and the non-tempral one as lst-fm.
6.
IAcontact [61]: This dataset represents a contact network among attendees of the ACM Hypertext conference in the year 2009. The links encode the interactions among the attendees during a 2.5-day period, while the nodes represent the attendees. In this paper, we refer the temporal version as iacontact-t and the non-temporal one as iacontact.
7.
FB-forum [61]: This dataset illustrates a forum-based online social network similar to Facebook. The network mostly records user actions and conversations in this forum. Individual nodes in this dataset represent users, and the links represent the forum messages these users have exchanged. The dataset spans a period of 164 days. In this paper, we refer the temporal version as forum-t and the non-temporal one as forum.
8.
Topology [62, 63]: This dataset encodes the complicated network of connections between autonomous systems on the Internet. Node refers to autonomous systems (AS), which are collections of connected IP routing domains run by independent network operators. In this paper, we refer the temporal version as topo-t and the non-temporal one as topo.
9.
Mooc [60]: This dataset displays user activity on a well-known MOOC program. The actions are symbolized as a temporal and directed network. Users are represented as nodes in the graph together with course activities (targets), and the links are user actions on the targets. In this paper, we refer the temporal version as act-mooc-t and the non-temporal one as act-mooc.
4.2 Networks without timestamps:
1.
ArXiv [59]: ArXiv is a network dataset representing collaborations among scientists who publish preprints and papers on the ArXiv platform. Nodes in the network represent individual authors, while links represent co-authorship connections. We refer the dataset prepared from this network as arxiv.
2.
Power-grid [61]: The Power-grid network dataset models the interconnected power grid infrastructure in the United States. It provides information on the transmission lines, nodes, and elements that make up the electrical system, enabling research into electricity distribution across different geographic areas. Nodes represent power plants, substations, transformers, and other infrastructure, while links depict the transmission lines connecting these elements. This dataset is valuable for studying the dynamics and efficiency of the U.S. electrical power grid. In this paper, we refer the dataset prepared from this network as pow-grid.
3.
Tech-routers [61]: It is a tech routers network, which represents the connections between routers in a technological infrastructure like the internet or computer networks. Nodes in the network represent individual routers, and links symbolize direct physical or logical links between matching routers. We refer the dataset prepared from this network as routers.
4.
Bio-yeast [61]: The Bio-yeast dataset is a biological network database focused on protein-protein interactions in the yeast species Saccharomyces cerevisiae. It represents proteins as nodes and shows connections between proteins based on physical binding, enzymatic processes, or co-expression patterns. We refer the dataset prepared from this network as bio-yeast.
5.
Facebook [64]: It is a social network, built on the Facebook platform, focuses on a single user (“ego”) and their connections with other users. Nodes represent Facebook users, and links represent friendships between them. We refer the dataset prepared from this network as fb.
6.
BlogCatalog [61]: This dataset is a social relationship network. The network is made up of bloggers and their friends and other social connections. Nodes in this network are represented as bloggers and the links represent social relationship among the bloggers. In this paper, we refer the dataset prepared from this network as blog-cat.
| Network | #Nodes | #Edges | Type | Time Span | Dataset | #Positives | #two-hop positives | #three-hop positives |
| CollegeMsg | 1899 | 20296 | Directed | 194 days | clg-msg-t | 2142 | 1068(49.80%) | 983(45.89%) |
| clg-msg | 510 | 335(65.68%) | 174(34.11%) | |||||
| Email-Eu | 986 | 332334 | Directed | 803 days | email-t | 3977 | 3774(94.89%) | 200(5.03%) |
| 441 | 429(97.27%) | 11(2.49%) | ||||||
| Mathoverflow | 24818 | 506550 | Directed | 2350 days | math-flow-t | 24225 | 18918(78.09%) | 4889(20.18%) |
| math-flow | 11373 | 9981(87.76%) | 1125(9.89%) | |||||
| Ast-Phys | 9756 | 86697 | Undirected | 3316 days | ast-phys-t | 32186 | 27035(83.99%) | 5151(16%) |
| ast-phys | 8898 | 8862(99.59%) | 36(0.40%) | |||||
| LastFm | 7624 | 27806 | Undirected | 1345 days | lst-fm-t | 29867 | 29862(99.38%) | 185(0.61%) |
| lst-fm | 3755 | 3748(99.81%) | – | |||||
| IAcontact | 113 | 2498 | Directed | 1 day | iacontact-t | 533 | 533(100%) | – |
| iacontact | 92 | 92(100%) | – | |||||
| FB-forum | 899 | 7089 | Directed | 164 days | forum-t | 1638 | 717(43.58%) | 862(52.81%) |
| forum | 277 | 163(58.84%) | 109(39.35%) | |||||
| Topology | 34761 | 171403 | Undirected | 23 days | topo-t | 21100 | 17331(82.13%) | 3159(14.97%) |
| topo | 4682 | 3578(76.42%) | 980(20.93%) | |||||
| Mooc | 7047 | 411749 | Directed | 29 days | act-mooc-t | 38448 | 38381(99.82%) | 67(0.01%) |
| act-mooc | 8594 | 8590(99.95%) | – | |||||
| ArXiv | 23133 | 93133 | Undirected | arxiv | 886 | 867(97.85%) | 6(0.14%) | |
| Power-grid | 4941 | 6594 | Undirected | pow-grid | 183 | 117(63.93%) | 65(35.51%) | |
| Tech-routers | 2113 | 6632 | Undirected | routers | 590 | 72(12.20%) | 45(7.62%) | |
| Bio-yeast | 6008 | 156945 | Undirected | bio-yeast | 1572 | 1472(93.63%) | 100(6.36%) | |
| 4039 | 88234 | Undirected | fb | 8813 | 8796(99.80%) | 17(0.19%) | ||
| BlogCatalog | 10312 | 333983 | Undirected | blog-cat | 33350 | 33112(99.28%) | 238(0.71%) |
5 Experimental Setup
5.1 Graph and ground truth generation
In this section, we detail the procedure of how we prepare the graph and the ground truth positive and negative node pairs from the datasets presented in Section 4, for future and missing LP tasks. Unlike traditional approaches, we carefully segregate the test examples based on the distance of the test node pairs in the graph, and maintain the distance bias inherent with the positive examples’ distribution, while sampling the negative examples. Moreover, it takes care of the evaluation in a class imbalance environment and its interactions with the factors stated above.
The future LP datasets were prepared for the networks where the interaction or link formation timestamps were available. The interactions can be directed, and there could be multiple interactions between a node pair. As we aim to prepare a simple and undirected graph and the true future links from a network, we ignored the directions of the interactions and the duplicate interactions. The detailed procedure for the dataset preparation is described as follows. We first sort the interactions or links (for the cases where only the link formation timestamps were available) by their timestamp values, and then choose of the earlier timestamped interactions or links for constructing the graph, and the remaining were used for generating positive examples. From the interactions which are in that , we constructed an unweighted and undirected graph by first ignoring the directions and then removing duplicate interactions and links, if any. If the graph is having multiple connected components, we choose the largest connected component as the final graph for our experiment. From the remaining interactions, we ignored the directions and removed the duplicates to generate the potential positive examples. From this potential set, we first ignored the node pairs which involve at least one node which is not included in the graph, and then removed the links which are present in the graph as an edge to get the final set of positive examples.
The missing LP was applied to all the networks presented in Section 4. We preprocessed the networks having timestamp information by ignoring the timestamps and their directions, and then removing the duplicates. Similarly, we ignored the directions and removed the duplicates, if applicable, for the networks which did not have timestamp information. After doing this, we are left with a set of node pairs which represents links. Then we randomly remove of the links from the set. We construct the graph with the remaining links and choose the largest connected component as the final graph. We consider the removed links as the potential positive examples, and generate the positive examples following the same procedure as the future LP datasets.
For each dataset, we divided the positive examples in three parts based on the shortest path distance of the end node pair in the graph: having the shortest path distance (a) , (b) , and (c) more than . As most of the datasets have no or a very few end node pairs having the shortest path distance , we do not consider this group in our experiment. To evaluate the effect of class imbalance on the LP methods, we generated negative examples times as many as the number of positive examples, stratified by specific shortest path distance categories maintaining the ratio of the representatives of the categories same as the positive set, and conducted experiments on multiple skew levels. For each category, we repeatedly choose two random nodes having the specified shortest path distance in the graph and if they are not included in the set of positive examples, we include the pair in the negative set. This sampling strategy empower us to evaluate the LP methods in groups of test examples by their shortest path distance in class imbalanced setup. The details of the positive example count for each dataset grouped by the distance categories are provided in Table 2.
5.2 Hyper-parameter settings for the LP methods
The Katz similarity method was implemented with , which is the typical value used in the literature [1]. Hyper-parameters of the node embedding models: Deepwalk [19], Nodevec [20], GraphSAGE [21], were set following the recommendations given in the actual paper. We performed five-fold cross validation, stratified by the skew level of the class distribution, to apply logistic regression and random forest on the link features towards predicting links.
5.3 Evaluation measures
The area under the receiver operating curve (AUROC) [31] has historically been the primary performance metric used to assess the performance of LP methods [2, 24]. The receiver operating curve (ROC) plots the true positive rate (TPR) against the false positive rate (FPR) at every threshold value. AUROC gives the probability of a positive example assigned a higher score than that of a negative example. However, AUROC is not a suitable performance measure in a class imbalance scenario like LP problem, which is inherently skewed towards the negative class. AUROC favors accurate classification of positive examples, at the cost of significantly misclassifying negative examples. Hence, AUROC can overestimate the performance of LP. An alternative to AUROC is area under the precision-recall curve (AUPR) [32, 3]. Precision-recall (PR) curve plots precision at every threshold value vs. its recall, i.e., TPR. As precision is sensitive to class imbalance unlike FPR, it can provide better evaluation of LP [32]. Despite AUPR’s suitability, it is less popular in LP literature, may be due to the variability of PR curves in different skew levels, which makes it difficult to compare the LP methods when multiple skewness is considered. To deal with class imbalance, another skew sensitive evaluation measure, is sometimes used to evaluate LP as a single threshold measure, where the positive example count is considered as a typical value of [65, 2].
In this work, we consider AUROC, AUPR and as single point summary metrics for evaluating LP methods. For the skew sensitive metrices: AUPR and , we consider both balanced as well as imbalanced (skew level ) test set. We choose the value to be the positive example count, and name it as . To understand the behavior of the LP methods in early retrieval stage, and the effect of class imbalance on it, we employ setting value to be the half of the positive example count, which we denote as . We also use the PR curve (PRC) along with it to understand the behavior covering all threshold values.
6 Results and Discussion
We applied the 58 LP methods presented in Section 3 on the 24 datasets defined in Section 4 in a controlled experimental setup proposed in Section 5. We evaluated their performance based on the four evaluation metrics: AUROC, AUPR, and on the test sets. As described in Section 5.1, we had different test sets depending on the shortest path distance between test node pairs, and each of these had two versions: balanced and imbalanced. For the skew sensitive metrics: AUPR, and , we consider evaluating both for balanced and imbalanced test set. We performed a series of statistical tests to test an array of carefully designed hypotheses in this controlled setup to interrogate how the LP methods perform in different situations depending on the shortest path distance between test node pairs, prediction type, network type, and class imbalance. We also investigate how the class imbalance affect the LP methods in early retrieval phase using proper statistical tests and PRC plots. We present the detailed results along with the insights acquired from them as follows.
6.1 Future LP Vs. Missing LP
In this section, we investigate whether the prediction performance of the LP methods differ on a particular network for the two cases: future LP and missing LP. Given a network, we apply paired t-test to compare the absolute values of the prediction performances of the prediction methods given a particular evaluation metric for the two cases. To find the level of the ordinal association between the two cases, we apply Kendall?s Tau Rank Correlation Coefficient (). We consider AUROC, and both balanced and imbalanced versions of AUPR and as the evaluation measure. We perform the experiment on the nine networks which had timestamps in distance controlled test set. Table 3 and Table 4 summarize the results of the two tests respectively.
Almost all paired t-test p-values are significant across distances and evaluation measures, and for almost all networks, the LP methods perform better for future LP than missing LP. The only exception is the two-hop case of LastFm network, where either the missing LP wins or there is no significant difference. When the test node pairs are generated randomly in a longitudinal network, the temporal relationship among the links are lost. The outperformance of future LP might be due to the adverse affect of this information loss in their missing LP counterparts. This also indicates that when an application demands future LP, but the ground truth links are generated in traditional way, i.e., removing a portion of links from the network randomly, the prediction performance is significantly underestimated, and vice versa. Table 4 shows that except for the Topology and Mooc network, all rank correlation values are extremely low, in few cases even negative. This indicates that, for a future LP, if the ground truth links are generated in traditional way, the prediction performance rankings of various LP methods vary significantly than if the ground truth links were generated with true future links. Our results indicate the necessity to evaluate a LP method in both setups and report its performance in both types of prediction tasks, in order to understand its effectiveness.
| NETWORK | AUROC | AUPR | Pr@P | ||||||||
| balanced | imbalanced | balanced | imbalanced | ||||||||
| P-Value | Winner | P-Value | Winner | P-Value | Winner | P-Value | Winner | P-Value | Winner | ||
| two-hop | CollegeMsg | .001*** | future | .001*** | future | .001*** | future | .001*** | future | .001*** | future |
| Email-Eu | .001*** | future | .001*** | future | .001*** | future | .001*** | future | .001*** | future | |
| Mathoverflow | .001*** | future | .001*** | future | .001*** | future | .001*** | future | .001*** | future | |
| Ast-phys | .001*** | future | .001*** | future | 0.003** | future | .001*** | future | 0.003** | future | |
| LastFm | .001*** | missing | 0.105 | — | 0.091 | — | .001*** | missing | .001*** | missing | |
| IAcontact | .001*** | future | .001*** | future | .001*** | future | .001*** | future | .001*** | future | |
| FB-forum | .001*** | future | .001*** | future | 0.052 | — | .001*** | future | .001*** | future | |
| Topology | .001*** | future | .001*** | future | .001*** | future | .001*** | future | .001*** | future | |
| Mooc | 0.011* | future | 0.018* | future | 0.004** | future | 0.006** | future | 0.004** | future | |
| three-hop | CollegeMsg | .001*** | future | .001*** | future | 0.943 | — | .001*** | future | 0.346 | — |
| Email-Eu | .001*** | future | .001*** | future | .001*** | future | .001*** | future | .001*** | future | |
| Mathoverflow | .001*** | future | .001*** | future | .001*** | future | .001*** | missing | .001*** | future | |
| Ast-phys | .001*** | future | .001*** | future | .001*** | future | 0.298 | — | .001*** | future | |
| LastFm | — | — | — | — | — | — | — | — | — | — | |
| IAcontact | — | — | — | — | — | — | — | — | — | — | |
| FB-forum | .001*** | future | .001*** | future | .001*** | future | .001*** | future | .001*** | future | |
| Topology | .001*** | future | .001*** | future | .001*** | future | 0.003** | future | .001*** | future | |
| Mooc | — | — | — | — | — | — | — | — | — | — | |
| NETWORK | AUROC | AUPR | Pr@P | |||
| balanced | imbalanced | balanced | imbalanced | |||
| two-hop | CollegeMsg | 0.184 | 0.184 | 0.219 | 0.156 | 0.203 |
| Email-Eu | 0.272 | 0.319 | 0.472 | 0.294 | 0.526 | |
| Mathoverflow | 0.339 | 0.386 | 0.414 | 0.326 | 0.488 | |
| Ast-phys | 0.487 | 0.433 | 0.377 | 0.525 | 0.393 | |
| lastFm | 0.477 | 0.465 | 0.381 | 0.462 | 0.335 | |
| IAcontact | 0.027 | 0.080 | 0.038 | 0.039 | 0.075 | |
| FB-forum | 0.321 | 0.347 | 0.438 | 0.311 | 0.271 | |
| Topology | 0.785 | 0.783 | 0.771 | 0.723 | 0.742 | |
| Mooc | 0.744 | 0.743 | 0.698 | 0.678 | 0.741 | |
| three-hop | CollegeMsg | 0.009 | 0.118 | 0.158 | 0.022 | 0.230 |
| Email-Eu | -0.030 | 0.085 | 0.167 | -0.144 | 0.181 | |
| Mathoverflow | -0.020 | -0.071 | 0.018 | 0.139 | 0.060 | |
| Ast-phys | 0.256 | 0.464 | 0.578 | 0.422 | 0.628 | |
| lastFm | — | — | — | — | — | |
| IAcontact | — | — | — | — | — | |
| FB-forum | 0.189 | 0.208 | 0.283 | 0.203 | 0.334 | |
| Topology | 0.749 | 0.747 | 0.738 | 0.649 | 0.773 | |
| Mooc | — | — | — | — | — | |
6.2 Distance between the end nodes in a test node pair
Here we aim to understand whether prediction performance of various LP methods vary when the distance between the two nodes in test node pairs differ. For each dataset, we performed paired t-test to determine if the prediction performance differ in terms of their absolute values for the two cases: two-hop away test node pairs and three-hop away test node pairs. Also, we measure the Kendall?s values for the the same to estimate the ordinal association between the two cases. We consider AUROC, AUPR (balanced), AUPR (imbalanced), (balanced) and (imbalanced) as the evaluation measures. We performed the experiment across all datasets and prediction types. We summarize the results of the two tests in Table 5 and Table 6 respectively.
Table 5 shows that the absolute performance of the prediction methods differs significantly in most of the cases across all datasets, prediction types and evaluation metrics. Overall, for a given dataset, the winner between two-hop and three-hop is consistent across the evaluation measures. As far as the missing LP is concerned, there is no clear winner between the two cases. However, for the future LP, three-hop seems to do better than its two-hop counterpart, where, three-hop wins in clg-msg-t, math-flow-t, ast-phys-t and lst-fm-t, but there is no clear winner for the rest three datasets. When we compare the performance for the networks (wherever applicable) based on prediction type, we find some inconsistencies, which is most prominent in the Ast-phys network, where its missing LP version is dominated by two-hop, but three-hop does better in its future LP counterpart. These results indicate that the LP methods produce different level of accuracy for predicting links based on the distance of the end nodes in the network. Moreover, this behavior differs over the prediction types. From Table 6 it is also evident that the rankings of the prediction methods for the two cases vary significantly. Overall, the obtained results support the need to evaluate LP methods separately for the two-hop and three-hop case to understand its behavior in distance controlled manner, and our experimental setup provides an opportunity to measure it. Moreover, as the local LP methods are applicable only on in the two-hop case, but the global methods are applicable in both the cases, our experimental setup provides a robust environment to compare a local with a global method.
| DATASET | AUROC | AUPR | Pr@P | ||||||||
| balanced | imbalanced | balanced | imbalanced | ||||||||
| P-Value | Winner | P-Value | Winner | P-Value | Winner | P-Value | Winner | P-Value | Winner | ||
| missing LP | clg-msg | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop |
| .001*** | 2-hop | .001*** | 2-hop | .001*** | 2-hop | .001*** | 2-hop | .001*** | 2-hop | ||
| math-flow | 0.226 | — | 0.223 | — | 0.492 | — | .001*** | 3-hop | 0.170 | — | |
| ast-phys | .001*** | 2-hop | .001*** | 2-hop | .001*** | 2-hop | .001*** | 3-hop | .001*** | 2-hop | |
| lst-fm | — | — | — | — | — | — | — | — | — | — | |
| forum | 0.007** | 2-hop | 0.134 | — | .001*** | 2-hop | 0.012* | 2-hop | 0.066 | — | |
| topo | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop | |
| act-mooc | — | — | — | — | — | — | — | — | — | — | |
| pow-grid | 0.512 | — | .001*** | 2-hop | .001*** | 2-hop | .001*** | 2-hop | .001*** | 2-hop | |
| routers | 0.923 | — | .001*** | 2-hop | 0.229 | — | .001*** | 2-hop | 0.029* | 3-hop | |
| bio-yeast | .001*** | 2-hop | .001*** | 2-hop | .001*** | 2-hop | .001*** | 3-hop | .001*** | 2-hop | |
| fb | .001*** | 3-hop | .001*** | 2-hop | 0.101 | — | .001*** | 3-hop | 0.682 | — | |
| blog-cat | .001*** | 2-hop | .001*** | 2-hop | .001*** | 2-hop | .001*** | 3-hop | .001*** | 2-hop | |
| future LP | clg-msg-t | .001*** | 3-hop | .008** | 3-hop | .007** | 3-hop | .005** | 3-hop | .004** | 3-hop |
| email-t | .001*** | 2-hop | 0.463 | — | 0.448 | — | .001*** | 2-hop | .049* | 3-hop | |
| math-flow-t | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop | |
| ast-phys-t | 0.081 | — | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop | |
| lst-fm-t | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop | .001*** | 3-hop | |
| forum-t | 0.827 | — | 0.255 | — | 0.531 | — | 0.563 | — | 0.212 | — | |
| topo-t | 0.334 | — | 0.585 | — | 0.004** | 3-hop | 0.017* | 2-hop | 0.013* | 3-hop | |
| act-mooc-t | 0.265 | — | .004** | 2-hop | 0.824 | — | .001*** | 3-hop | 0.372 | — | |
| DATASET | AUROC | AUPR | Pr@P | |||
| balanced | imbalanced | balanced | imbalanced | |||
| missing LP | clg-msg | 0.275 | 0.203 | 0.241 | 0.316 | 0.207 |
| -0.059 | -0.012 | -0.155 | -0.091 | 0.007 | ||
| math-flow | 0.399 | 0.385 | 0.227 | 0.311 | 0.196 | |
| ast-phys | 0.278 | 0.391 | 0.572 | 0.373 | 0.542 | |
| lst-fm | — | — | — | — | — | |
| forum | 0.184 | 0.227 | 0.236 | 0.221 | 0.221 | |
| topo | 0.233 | 0.316 | 0.302 | 0.176 | 0.241 | |
| act-mooc | — | — | — | — | — | |
| pow-grid | -0.144 | 0.084 | 0.173 | -0.108 | 0.189 | |
| routers | 0.242 | 0.005 | 0.079 | 0.285 | 0.068 | |
| bio-yeast | 0.216 | 0.260 | 0.362 | 0.316 | 0.302 | |
| fb | -0.178 | 0.450 | 0.211 | -0.159 | 0.025 | |
| blog-cat | 0.213 | 0.031 | 0.049 | 0.319 | 0.033 | |
| future LP | clg-msg-t | 0.580 | 0.589 | 0.548 | 0.578 | 0.489 |
| email-t | 0.603 | 0.443 | 0.390 | 0.609 | 0.337 | |
| math-flow-t | 0.387 | 0.437 | 0.467 | 0.391 | 0.445 | |
| ast-phys-t | 0.660 | 0.705 | 0.677 | 0.508 | 0.653 | |
| lst-fm-t | 0.112 | -0.021 | -0.095 | 0.222 | -0.131 | |
| forum-t | 0.476 | 0.467 | 0.511 | 0.488 | 0.406 | |
| topo-t | 0.363 | 0.391 | 0.360 | 0.326 | 0.361 | |
| act-mooc-t | 0.439 | 0.589 | 0.573 | 0.413 | 0.594 | |
6.3 Prediction method type
Here we investigate if the prediction performance differ with the type of the LP methods. As all of the local-sim, global-sim and learning based methods apply on the two-hop case, we perform one way ANOVA followed by Tukey post-hoc test on these three groups of methods. As only global-sim and learning based methods apply on the three-hop case, we apply Student’s t-test for this case. We summarize the results of the ANOVA for two-hop in Table 7, Tukey test results for two-hop in Table 8 and Table 9 respectively for missing LP and future LP, and the t-test results for the three-hop case in Table 10.
Table 7 shows that the prediction performance of the prediction methods in the three groups differ significantly for almost all cases in two-hop. Post-hoc test for the missing LP in two-hop cases presented in Table 8 shows that across almost all datasets learning group is the clear winner, but although global-sim methods outperforms local-sim in a few datasets and evaluation methods, in most cases there are no significant difference. For the two-hop cases in future LP shown in Table 9, although in clg-msg-t, math-flow-t and math-flow-t datasets, learning methods perform best, results are not conclusive for the rest. Table 10 shows that, for three-hop case, there is no significant difference in performance between the global-sim and learning based methods, except for few cases like bio-yeast, forum, blog-cat and lst-fm-t. For bio-yeast, forum and blog-cat, global-sim methods win, and learning based methods win in lst-fm-t. However, the number of the test node pairs for bio-yeast, forum and blog-cat in three-hop case is low, so, over-fitting might have caused the performance degradation for the learning based methods. Overall, the results indicate that the relative performance of different types of LP methods vary with the distance between the end nodes. Moreover, even when the distance is same, the relative performance of the method types vary with the prediction type. The results in this section advocate for the proposed distance controlled and prediction type controlled setup, along with the consideration of the type of the methods while comparing performance of multiple methods, in order to evaluate an LP method in a robust and detailed manner.
| DATASET | AUROC | AUPR | Pr@P | |||
| balanced | imbalanced | balanced | imbalanced | |||
| missing LP | clg-msg | .001*** | .001*** | .001*** | .001*** | .001*** |
| .001*** | .001*** | .001*** | .001*** | .001*** | ||
| math-flow | .001*** | .001*** | .001*** | .001*** | .001*** | |
| ast-phys | 0.014* | 0.016** | 0.001** | 0.019* | 0.015** | |
| lst-fm | .001*** | .001*** | .001*** | .001*** | .001*** | |
| iacontact | 0.136 | 0.126 | .001*** | 0.508 | 0.012** | |
| forum | .001*** | .001*** | .001*** | .001*** | .001*** | |
| topo | 0.005** | .001*** | .001*** | 0.028* | .001*** | |
| act-mooc | 0.541 | 0.449 | 0.437 | 0.687 | 0.410 | |
| arxiv | 0.001** | .001*** | .001*** | .001*** | .001*** | |
| pow-grid | 0.415 | 0.003** | .001*** | 0.012* | 0.008** | |
| routers | .001*** | .001*** | .001*** | .001*** | .001*** | |
| bio-yeast | .001*** | .001*** | .001*** | .001*** | .001*** | |
| fb | 0.070 | 0.011* | .001*** | 0.021* | 0.003** | |
| blog-cat | .001*** | .001*** | .001*** | .001*** | .001*** | |
| future LP | clg-msg-t | .001*** | .001*** | .001*** | .001*** | .001*** |
| email-t | 0.161 | 0.012* | .001*** | 0.157 | .001*** | |
| math-flow-t | .001*** | .001*** | .001*** | .001*** | .001*** | |
| ast-phys-t | 0.007** | .001*** | 0.003** | 0.104 | .001*** | |
| lst-fm-t | 0.001** | 0.002** | 0.001** | .001*** | 0.002** | |
| iacontact-t | 0.096 | 0.008** | .001*** | 0.162 | .001*** | |
| forum-t | .001*** | 0.003** | .001*** | .001*** | 0.027* | |
| topo-t | 0.085 | 0.042* | .001*** | 0.046* | .001*** | |
| act-mooc-t | .001*** | .001*** | .001*** | .001*** | .001*** | |
DATASET METHOD GROUP AUROC AUPR Pr@P balanced imbalanced balanced imbalanced P-Value Winner P-Value Winner P-Value Winner P-Value Winner P-Value Winner clg-msg learning vs global-sim .001*** learning .001*** learning .001*** learning .001*** learning .001*** learning global-sim vs local-sim 0.762 — 0.787 — 0.631 — 0.590 — 0.415 — learning vs local-sim .001*** learning .001*** learning .001*** learning .001*** learning .001*** learning email learning vs global-sim 0.006** learning 0.004** learning 0.001** learning 0.001** learning 0.001** learning global-sim vs local-sim 0.086 — 0.022* global-sim 0.031* global-sim .001*** global-sim 0.006** global-sim learning vs local-sim .001*** learning .001*** learning .001*** learning .001*** learning .001*** learning math-flow learning vs global-sim .001*** learning .001*** learning .001*** learning .001*** learning .001*** learning global-sim vs local-sim 0.459 — 0.518 — 0.297 — 0.116 — 0.353 — learning vs local-sim .001*** learning .001*** learning .001*** learning .001*** learning .001*** learning ast-phys learning vs global-sim 0.221 — 0.251 — 0.181 — 0.049* learning 0.178 — global-sim vs local-sim 0.767 — 0.738 — 0.894 — 0.804 — 0.862 — learning vs local-sim 0.029* learning 0.030* learning 0.050* learning 0.005** learning 0.039** learning lst-fm learning vs global-sim .001*** learning 0.007** learning .001*** learning 0.005** learning 0.572 — global-sim vs local-sim 0.267 — 0.045* global-sim 0.296 — 0.041* global-sim 0.020** global-sim learning vs local-sim .001*** learning .001*** learning .001*** learning .001*** learning .001*** learning iacontact learning vs global-sim 0.452 — 0.606 — 0.507 — 0.376 — 0.924 — global-sim vs local-sim 0.117 — 0.123 — 0.568 — 0.008** global-sim 0.040** global-sim learning vs local-sim 0.272 — 0.186 — 0.958 — .001*** learning 0.010** learning forum learning vs global-sim .001*** learning .001*** learning .001*** learning .001*** learning .001*** learning global-sim vs local-sim 0.217 — 0.895 — 0.208 — 0.630 — 0.674 — learning vs local-sim 0.007** learning .001*** learning 0.038* learning .001*** learning .001*** learning topo learning vs global-sim 0.197 — 0.050* learning 0.481 — 0.001** learning 0.005** learning global-sim vs local-sim 0.602 — 0.656 — 0.558 — 0.188 — 0.472 — learning vs local-sim 0.009** learning 0.002** learning 0.0336* learning .001*** learning .001*** learning act-mooc learning vs global-sim 0.624 — 0.545 — 0.876 — 0.878 — 0.830 — global-sim vs local-sim 0.980 — 0.977 — 0.981 — 0.851 — 0.879 — learning vs local-sim 0.782 — 0.721 — 0.738 — 0.441 — 0.432 — arxiv learning vs global-sim 0.221 — 0.251 — 0.181 — 0.999 — 0.980 — global-sim vs local-sim 0.767 — 0.738 — 0.894 — .001*** global-sim 0.001** global-sim learning vs local-sim 0.029* learning 0.030* learning 0.050* learning .001*** learning .001*** learning pow-grid learning vs global-sim 0.562 — 0.148 — 0.930 — 0.040* learning 0.015** learning global-sim vs local-sim 0.996 — 0.626 — 0.041* global-sim 0.004** global-sim 0.666 — learning vs local-sim 0.635 — 0.007** learning 0.009** learning .001*** learning 0.212 — routers learning vs global-sim 0.059 — 0.008** learning 0.057 — .001*** learning .001*** learning global-sim vs local-sim 0.331 — 0.054 — 0.707 — .001*** global-sim 0.002** global-sim learning vs local-sim .001*** learning .001*** learning 0.003** learning .001*** learning .001*** learning bio-yeast learning vs global-sim 0.004** learning 0.017* learning 0.002** learning 0.004** learning 0.006** learning global-sim vs local-sim 0.692 — 0.600 — 0.494 — 0.130 — 0.194 — learning vs local-sim .001*** learning .001*** learning .001*** learning .001*** learning .001*** learning fb learning vs global-sim 0.657 — 0.414 — 0.633 — 0.173 — 0.335 — global-sim vs local-sim 0.582 — 0.456 — 0.368 — 0.167 — 0.334 — learning vs local-sim 0.073 — 0.014* learning 0.020* learning .001*** learning 0.004** learning blog-cat learning vs global-sim 0.002** learning 0.001** learning 0.001** learning .001*** learning .001*** learning global-sim vs local-sim 0.774 — 0.816 — 0.680 — 0.546 — 0.712 — learning vs local-sim .001*** learning .001*** learning .001*** learning .001*** learning .001*** learning
DATASET METHOD GROUP AUROC AUPR Pr@P balanced imbalanced balanced imbalanced P-Value Winner P-Value Winner P-Value Winner P-Value Winner P-Value Winner clg-msg-t learning vs global-sim 0.001** learning 0.007** learning .001*** learning .001*** learning 0.001** learning global-sim vs local-sim 0.366 — 0.933 — 0.785 — 0.109 — 0.706 — learning vs local-sim 0.159 — 0.026* learning 0.002** learning 0.299 — 0.027* learning email-t learning vs global-sim 0.970 — 0.879 — 0.715 — 0.977 — 0.801 — global-sim vs local-sim 0.432 — 0.149 — .001*** global-sim 0.414 — 0.004** global-sim learning vs local-sim 0.137 — 0.009** learning .001*** learning 0.133 — .001*** learning math-flow-t learning vs global-sim 0.001** learning .001*** learning .001*** learning .001*** learning .001*** learning global-sim vs local-sim 0.830 — 0.764 — 0.204 — 0.787 — 0.521 — learning vs local-sim .001*** learning .001*** learning .001*** learning .001*** learning .001*** learning ast-phys-t learning vs global-sim 0.037* global-sim 0.004** global-sim 0.072 — 0.173 — 0.011* global-sim global-sim vs local-sim 0.974 — 1.000 — 0.873 — 0.922 — 0.615 — learning vs local-sim 0.074 — 0.005** local-sim 0.014* local-sim 0.404 — .001*** local-sim lst-fm-t learning vs global-sim 0.852 — 0.772 — 0.957 — 0.433 — 0.083 — global-sim vs local-sim 0.029* global-sim 0.008** global-sim 0.010* global-sim 0.063 — 0.002** global-sim learning vs local-sim .001*** learning 0.002** learning .001*** learning .001*** learning 0.015* learning iacontact-t learning vs global-sim 0.911 — 0.568 — 0.928 — 0.935 — 0.222 — global-sim vs local-sim 0.159 — 0.013* global-sim .001*** global-sim 0.487 — .001*** global-sim learning vs local-sim 0.090 — 0.011* learning .001*** learning 0.141 — .001*** learning forum-t learning vs global-sim 0.001** learning 0.018* learning 0.006** learning .001*** learning 0.058 — global-sim vs local-sim 0.265 — 0.962 — 0.992 — 0.003** local-sim 0.858 — learning vs local-sim 0.167 — 0.045* learning 0.009** learning 0.744 — 0.244 — topo-t learning vs global-sim 0.660 — 0.516 — 0.200 — 0.521 — 0.157 — global-sim vs local-sim 0.625 — 0.607 — 0.277 — 0.622 — 0.366 — learning vs local-sim 0.091 — 0.050* learning .001*** learning 0.055 — .001*** learning act-mooc-t learning vs global-sim 0.003** global-sim .001*** global-sim 0.086 — 0.024* global-sim 0.079 — global-sim vs local-sim 0.939 — 0.859 — 0.583 — 0.825 — 0.468 — learning vs local-sim 0.001** local-sim .001*** local-sim 0.003** local-sim 0.003** local-sim .001*** local-sim
DATASET AUROC AUPR Pr@P balanced imbalanced balanced imbalanced P-Value Winner P-Value Winner P-Value Winner P-Value Winner P-Value Winner missing LP clg-msg 0.043* global-sim 0.054 — 0.082 — 0.043* global-sim 0.057 — email 0.039* global-sim 0.157 — 0.078 — .001*** global-sim 0.532 — math-flow 0.365 — 0.206 — 0.195 — 0.462 — 0.199 — ast-phys 0.329 — 0.479 — 0.303 — 0.232 — 0.388 — lst-fm — — — — — — — — — — iacontact — — — — — — — — — — forum 0.020* global-sim 0.039* global-sim 0.048* global-sim 0.016* global-sim 0.036* global-sim topo 0.558 — 0.173 — 0.215 — 0.694 — 0.662 — act-mooc — — — — — — — — — — arxiv — — — — — — — — — — pow-grid 0.367 — 0.287 — 0.164 — 0.581 — 0.251 — routers 0.458 — 0.325 — 0.470 — 0.391 — 0.463 — bio-yeast .001*** global-sim 0.019* global-sim 0.012* global-sim .001*** global-sim 0.003* global-sim fb 0.328 — 0.650 — 0.552 — .001*** learning 0.925 — blog-cat 0.007** global-sim 0.004** global-sim 0.003** global-sim 0.094 — .001*** global-sim future LP clg-msg-t 0.288 — 0.328 — 0.227 — 0.293 — 0.150 — email-t 0.611 — 0.599 — 0.165 — 0.400 — 0.241 — math-flow-t 0.516 — 0.356 — 0.297 — 0.528 — 0.287 — ast-phys-t 0.705 — 0.554 — 0.527 — 0.730 — 0.545 — lst-fm 0.017* learning 0.473 — 0.256 — 0.010** learning 0.034* learning iacontact — — — — — — — — — — forum-t 0.051 — 0.052 — 0.117 — 0.044* global-sim 0.073 — topo-t 0.616 — 0.248 — 0.108 — 0.606 — 0.337 — act-mooc-t 0.613 — 0.717 — 0.898 — 0.785 — 0.998 —
6.4 Network type
As most of the LP methods apply on undirected networks, while applying those methods on directed networks, they are converted to undirected ones ignoring directions. Here we investigate whether this information loss affects their prediction performance. For each of future and missing LP, we divide the datasets into two groups: directed and undirected, and perform Student’s t-test for a LP method given its accuracy against a particular evaluation metric. To avoid getting overloaded with the high number of the LP methods, we conduct the test once for each type of methods: local-sim, global-sim and learning, by considering the performance score of the best performing method in that group. In Table 11, we report the test results. For the best performing methods in each group given an evaluation metric, readers may refer to Appendix A.
We find few tests achieve the significance level, and for all such cases the undirected networks are the winner. More interestingly, the undirected version wins for all cases for the future LP when the evaluation method is the imbalanced version of AUPR or . This evaluation setup is the most prominent: we are considering future LP on a longitudinal network, and our test set is imbalanced. The results in this particular setup is indicative for the hypothesis that the information loss in directed networks due to ignoring directions may adversely affect the prediction performance. Note that, we do not strike out the possibility of the presence of any confounder inherent to the directed networks which may result the performance degradation. Study on such confounders is out of the scope of this paper.
DATASET METHOD TYPE AUROC AUPR Pr@P balanced imbalanced balanced imbalanced P-Value Winner P-Value Winner P-Value Winner P-Value Winner P-Value Winner future two-hop local-sim 0.736 — 0.880 — 0.025* undirected 0.687 — 0.018* undirected global-sim 0.750 — 0.787 — 0.036* undirected 0.770 — 0.045* undirected learning 0.737 — 0.688 — 0.038* undirected 0.681 — 0.045* undirected missing two-hop local-sim 0.199 — 0.091 — 0.030* undirected 0.152 — 0.072 — global-sim 0.577 — 0.433 — 0.607 — 0.512 — 0.590 — learning 0.384 — 0.322 — 0.566 — 0.432 — 0.541 — future three-hop global-sim 0.716 — 0.757 — 0.593 — 0.049* undirected 0.593 — learning 0.343 — 0.376 — 0.416 — 0.068 — 0.416 — missing three-hop global-sim 0.505 — 0.381 — 0.119 — 0.197 — 0.114 — learning 0.020* undirected 0.062 — 0.062 — 0.173 — 0.093 —
6.5 Class imbalance and Early retrieval
In this section, we analyze how class imbalance affects the evaluation of LP. First we investigate if the relative performance of the LP methods differ subject to AUROC vs. AUPR. Both of AUROC and AUPR consider all threshold values, but AUROC is not affected by class imbalance whereas AUPR is. We consider both balanced as well as imbalanced (skew level ) test set for AUPR. Next, we take the imbalance sensitive measure AUPR, and compare the relative performance of the LP methods on it for the balanced and imbalanced version of the test set. We further do the same with the single threshold imbalance sensitive metric . We consider Kendall?s Tau Rank Correlation Coefficient () to measure the (in)consistently of ranks of the LP methods when evaluated in each pair of setups stated above. The values for each of the datasets are shown in Table 12. We summarize our observations as follows.
Considering the correlation values for AUROC vs. AUPR, it is evident that for almost every dataset across all distances (two-hop/three-hop) and prediction types (future LP/missing LP), the correlation decreases when the imbalanced version of the test set is considered over the balanced one. For a significant number of cases (such as, in two-hop scenario, iacontact, forum, arxiv, pow-grid, etc.) the correlation in imbalance scenario are considerably low. This indicates that when tested on imbalance dataset, there are significant disagreements on the performance of various LP methods when AUPR is used against AUROC. The correlation values for AUPR: balanced vs. imbalanced, and : balanced vs. imbalanced indicate some disagreements, which is higher for . This observation is consistent across distances and prediction types. These results show that if LP methods are tested on balanced test set, the conclusions may be erroneous, because they may behave differently in an imbalance setup which is inherent to the LP problem. It also suggests that although AUROC has been traditionally used to evaluate LP methods, their performance significantly vary if skew sensitive metrics like AUPR or are applied on an imbalanced test setup. Hence we recommend AUPR or as a single point summary metric applied on an imbalanced setup to evaluate LP, while maintaining other factor as described in Section 5.1. Unlike , as AUPR covers all threshold values, it can capture the whole spectrum of the precision-recall space, it should be the first choice.
6.5.1 Early Retrieval:
For certain applications like recommender systems, LP methods must perform well at the early retrieval phase, and it becomes critical when the data is highly skewed. Here we investigate how class imbalance affects early retrieval. The ROC curve, rather the Concentrated ROC (CROC) curve [34] can reflect the performance in the early retrieval stage. However, as it is skew insensitive, it is unable to reflect the performance in imbalanced setup. Saito et al. [32] has shown how PRC can be used to understand the effect of imbalance in early retrieval. We first investigate how the performance on the single point metric (both balanced and imbalanced versions) correlate with that of AUROC, and the balanced and imbalanced versions of correlate with each other, and then show an example ROC plot demonstrating the different behavior of the prediction methods in the two test settings. For each dataset, we measure the Kendall?s Tau Rank Correlation Coefficient () of the LP methods for the pair of cases stated above in our controlled setup.
Table 13 shows that the correlations with AUROC are significantly decreasing from -balanced to -imbalanced case, which implies that with increasing imbalance, AUROC becomes very inconsistent with the early retrieval behavior. In many cases, the correlation between the balanced and the imbalanced version of is considerably low, which indicates that, in early retrieval phase if we consider imbalanced test set, LP methods behave differently as that of the balanced case. So, we recommend carefully investigating the early retrieval performance of the prediction methods considering imbalanced test set in the controlled setup presented in this paper, using proper visualization technique like PR curve. Figure 1 presents a case where the imbalance affect the early retrieval. It shows the PRC plots for the two methods and DW_L1_LR for the email dataset in two-hop test scenario, both for balanced and imbalanced cases. It shows that although they perform comparably in both of the balanced and imbalanced cases, outperforms DW_L1_LR clearly in the early retrieval phase in the imbalanced case.
DATASET two-hop three-hop AUROC vs AUROC vs AUPR (balanced) vs Pr@P (balanced) vs AUROC vs AUROC vs AUPR (balanced) vs Pr@P (balanced) vs AUPR (balanced) AUPR (imbalanced) AUPR (imbalanced) Pr@P (imbalanced) AUPR (balanced) AUPR (imbalanced) AUPR (imbalanced) Pr@P (imbalanced) missing LP clg-msg 0.81 0.72 0.75 0.54 0.73 0.46 0.58 0.39 email 0.86 0.75 0.76 0.56 — — — — math-flow 0.90 0.79 0.87 0.71 0.87 0.78 0.85 0.53 ast-phys 0.89 0.84 0.93 0.85 0.17 0.17 0.66 0.47 lst-fm 0.81 0.75 0.84 0.61 — — — — iacontact 0.73 0.48 0.56 0.21 — — — — forum 0.80 0.65 0.72 0.45 0.81 0.67 0.70 0.37 topo 0.85 0.81 0.95 0.64 0.79 0.59 0.72 0.58 act-mooc 0.95 0.91 0.95 0.87 — — — — arxiv 0.77 0.54 0.71 0.56 — — — — pow-grid 0.65 0.52 0.71 -0.14 0.79 0.56 0.67 0.30 routers 0.74 0.71 0.80 0.51 0.77 0.55 0.67 0.51 bio-yeast 0.86 0.78 0.89 0.77 0.49 0.38 0.59 0.09 fb 0.90 0.80 0.87 0.73 — — — — blog-cat 0.89 0.83 0.91 0.83 0.70 0.60 0.64 0.45 future LP clg-msg-t 0.84 0.82 0.84 0.66 0.85 0.84 0.86 0.73 email-t 0.76 0.67 0.82 0.51 0.87 0.71 0.73 0.59 math-flow-t 0.93 0.88 0.94 0.83 0.91 0.84 0.90 0.84 ast-phys-t 0.90 0.77 0.87 0.78 0.86 0.83 0.85 0.58 lst-fm-t 0.82 0.80 0.93 0.66 0.35 0.58 0.68 0.16 iacontact-t 0.82 0.80 0.78 0.56 — — — — forum-t 0.83 0.79 0.75 0.54 0.89 0.84 0.84 0.72 topo-t 0.91 0.85 0.93 0.85 0.91 0.76 0.82 0.76 act-mooc-t 0.93 0.87 0.93 0.87 0.57 0.74 0.81 0.39
DATASET two-hop three-hop AUROC vs Pr@P/2 (balanced) AUROC vs Pr@P/2 (imbalanced) Pr@P/2 (balanced) vs Pr@P/2 (imbalanced) AUROC vs Pr@P/2 (balanced) AUROC vs Pr@P/2 (imbalanced) Pr@P/2 (balanced) vs Pr@P/2 (imbalanced) missing LP clg-msg 0.769 0.514 0.550 0.618 0.369 0.413 email 0.769 0.460 0.380 0.713 0.107 0.185 math-flow 0.836 0.691 0.737 0.874 0.727 0.791 ast-phys 0.840 0.786 0.890 0.571 0.121 0.048 lst-fm 0.736 0.598 0.677 — — — iacontact 0.619 0.212 0.140 — — — forum 0.749 0.566 0.449 0.627 0.477 0.449 topo 0.799 0.776 0.862 0.533 0.455 0.784 act-mooc 0.826 0.761 0.898 — — — arxiv 0.685 0.552 0.692 — — — pow-grid 0.563 -0.009 -0.080 0.622 0.335 0.358 routers 0.812 0.133 0.131 0.717 0.243 0.197 bio-yeast 0.804 0.740 0.830 0.717 0.302 0.387 fb 0.822 0.687 0.721 0.271 0.163 0.181 blog-cat 0.857 0.775 0.877 0.763 0.525 0.503 future LP clg-msg-t 0.815 0.601 0.691 0.817 0.669 0.669 email-t 0.701 0.453 0.604 0.751 0.521 0.601 math-flow-t 0.921 0.801 0.862 0.867 0.773 0.845 ast-phys-t 0.802 0.659 0.824 0.564 0.598 0.703 lst-fm-t 0.798 0.625 0.677 0.642 0.312 0.319 iacontact-t 0.725 0.501 0.461 — — — forum-t 0.785 0.576 0.568 0.830 0.685 0.708 topo-t 0.866 0.795 0.892 0.738 0.623 0.763 act-mooc-t 0.839 0.788 0.913 0.280 0.302 0.161
6.6 Summary
Our proposed experimental setup allowed us to conduct the tests presented in this section in order to evaluate the existing LP methods in a controlled environment. Our main findings are: (a) LP methods perform better when future LP is considered than that of missing LP; (b) behavior of the LP methods are not consistent on the two test sets: two-hop and three-hop away node pairs; (c) when absolute performance of the LP methods are compared based on their category, we found significant difference in performance for the test set consisting of two-hop away node pairs, but for its three-hop away counterpart, no significant difference was observed; (d) converting directed networks in undirected ones and applying LP methods on it may significantly affect its performance for future LP on the test set of two-hop away node pairs; (e) injecting imbalance in the test set affects the relative performance of the LP methods for the skew sensitive metrics like AUPR and , and the rankings based on AUPR and are not consistent with that of the skew insensitive metric AUROC; (f) rankings of the LP methods differ based on their performance in the early retrieval phase when imbalanced test set used as opposed to balanced one.
Based on our findings, we recommend evaluating the LP methods in the controlled experimental setup presented in this paper. The methods should be evaluated on both the problem types, on the carefully designed distance controlled and skew controlled test set. A method’s performance needs to be compared within its category in this setup, as well as with all the methods irrespective of their categories. We recommend AUPR as a single point summary evaluation measure to be employed on both of the balanced and imbalanced test set. Along with it, PRC, on both of balanced and imbalanced test set, should be investigated to focus on certain portion of the whole spectrum of the thresholds, such as the early retrieval region, which may be necessary for certain applications.
7 Conclusion and Future Directions
In this paper, we identified a number of gaps in the existing link prediction literature concerning its performance evaluation. This paper proposed an experimental setup which provides a controlled environment based on factors like prediction type, network type, method type, distance between end nodes, class imbalance, evaluation metric, early retrieval scenario, etc., to evaluate LP methods in a rigorous manner. We performed extensive experiments on real network datasets using existing LP methods, and tested carefully designed hypotheses to gain insight on the effect of those factors on LP methods. Subsequently, we presented our recommendations to be followed as best practice toward effective evaluation of LP methods.
This paper considered evaluating link prediction on a simple, unweighted, undirected and homogeneous graph, and analyzed the LP methods which ignore time information. This work can be extended to identify the unexplored avenues in evaluation of LP on specialized graphs, such as weighted, directed and heterogeneous. Various aspects of the performance of temporal LP methods and LP methods, which are specifically designed for the directed networks, may be investigated as further studies.
8 Declarations
8.1 Ethical Approval
This declaration is “not applicable”.
8.2 Funding
This declaration is “not applicable”.
Appendix A Consolidated results: Best performing methods in individual groups of methods
DATASET METHOD TYPE AUROC AUPR Pr@P balanced imbalanced balanced imbalanced Method Name Value Method Name Value Method Name Value Method Name Value Method Name Value clg-msg-t local-sim RA 0.678 RA 0.685 RA 0.206 RA 0.630 RA 0.245 global-sim PA 0.751 PA 0.751 PA 0.267 PA 0.693 PA 0.300 learning N2V_Avg_LR 0.688 N2V_Avg_LR 0.688 DW_L2_LR 0.189 N2V_Avg_LR 0.643 N2V_Avg_LR 0.236 email-t local-sim RA 0.860 RA 0.846 RA 0.432 RA 0.786 RA 0.464 global-sim Katz 0.777 Katz 0.753 Katz 0.270 Katz 0.716 Katz 0.310 learning GS_MaxPool_Hada_RF 0.803 DW_L2_LR 0.764 DW_L2_LR 0.278 GS_MaxPool_Hada_RF 0.742 DW_L1_LR 0.348 math-flow-t local-sim RA 0.889 RA 0.902 RA 0.624 RA 0.811 RA 0.589 global-sim PA 0.912 PA 0.911 PA 0.621 PA 0.829 PA 0.579 learning DW_Avg_LR 0.871 DW_Avg_LR 0.860 DW_Avg_LR 0.438 DW_Avg_LR 0.793 DW_Avg_LR 0.465 ast-phys-t local-sim CN 0.789 RA 0.716 RA 0.196 AA 0.721 RA 0.184 global-sim PA 0.941 PA 0.908 PA 0.485 PA 0.856 PA 0.580 learning N2V_Avg_LR 0.634 N2V_Avg_RF 0.974 N2V_Avg_LR 0.756 N2V_Avg_RF 0.945 N2V_Avg_RF 0.736 lst-fm-t local-sim RA 0.591 RA 0.578 RA 0.126 RA 0.565 RA 0.151 global-sim PA 0.598 PA 0.581 PA 0.127 PA 0.570 PA 0.149 learning N2V_Avg_LR 0.576 N2V_Avg_LR 0.565 GS_Mean_Avg_RF 0.120 N2V_Avg_LR 0.557 GS_Mean_Avg_LR 0.148 iacontact-t local-sim RA 0.608 RA 0.608 RA 0.153 AA 0.573 RA 0.180 global-sim PA 0.583 PA 0.584 PA 0.133 PA 0.575 PA 0.167 learning GS_LSTM_L1_RF 0.612 GS_LSTM_L1_RF 0.614 GS_LSTM_L1_RF 0.134 GS_LSTM_L1_RF 0.618 GS_Mean_L2_RF 0.146 forum-t local-sim AA 0.608 AA 0.606 AA 0.146 AA 0.565 AA 0.179 global-sim PA 0.710 PA 0.673 PA 0.193 PA 0.642 PA 0.241 learning DW_Avg_LR 0.641 DW_Avg_RF 0.619 GS_MeanPool_Avg_RF 0.151 DW_Avg_LR 0.596 DW_Avg_RF 0.197 topo-t local-sim AA 0.938 AA 0.945 AA 0.609 AA 0.857 AA 0.581 global-sim PA 0.952 PA 0.951 PA 0.617 PA 0.890 PA 0.588 learning DW_L2_LR 0.914 DW_Avg_LR 0.899 GS_MaxPool_Hada_RF 0.552 DW_Avg_LR 0.832 DW_L1_LR 0.570 act-mooc-t local-sim RA 0.738 RA 0.713 RA 0.23 CN 0.660 RA 0.237 global-sim PA 0.974 PA 0.963 PA 0.763 PA 0.945 PA 0.836 learning DW_Avg_LR 0.977 N2V_Hada_RF 0.976 N2V_Hada_LR 0.857 N2V_Avg_RF 0.945 N2V_L1_LR 0.833
DATASET METHOD TYPE AUROC AUPR Pr@P balanced imbalanced balanced imbalanced Method Name Value Method Name Value Method Name Value Method Name Value Method Name Value clg-msg local-sim RA 0.721 RA 0.720 RA 0.222 RA 0.687 RA 0.281 global-sim Katz 0.799 Katz 0.805 Katz 0.280 Katz 0.746 Katz 0.356 learning DW_L2_LR 0.661 DW_L2_LR 0.649 DW_L2_LR 0.171 DW_L2_LR 0.643 DW_L1_LR 0.201 email local-sim RA 0.855 RA 0.842 RA 0.325 RA 0.807 RA 0.394 global-sim Katz 0.791 Katz 0.775 Katz 0.257 Katz 0.744 Katz 0.322 learning DW_L2_LR 0.776 DW_L2_LR 0.759 DW_L2_LR 0.231 DW_L2_LR 0.715 DW_L2_LR 0.278 math-flow local-sim AA 0.878 AA 0.885 AA 0.545 AA 0.808 AA 0.546 global-sim PA 0.884 PA 0.879 PA 0.511 KPA 0.800 Katz 0.529 learning DW_L2_LR 0.953 DW_L2_LR 0.809 DW_L2_LR 0.380 DW_L2_LR 0.729 DW_L2_LR 0.437 ast-phys local-sim RA 0.920 RA 0.952 RA 0.661 RA 0.900 RA 0.631 global-sim PA 0.984 CT 0.828 PA 0.826 CT 0.793 PA 0.820 learning N2V_L2_LR 0.953 GS_Mean_L1_LR 0.819 N2V_L1_LR 0.620 GS_MeanPool_L2_LR 0.745 N2V_L2_LR 0.650 lst-fm local-sim RA 0.667 RA 0.621 AA 0.147 RA 0.617 CN 0.178 global-sim PA 0.675 PA 0.626 PA 0.150 PA 0.622 PA 0.177 learning DW_L2_LR 0.623 DW_L2_LR 0.577 DW_L2_LR 0.125 DW_L2_LR 0.585 DW_L1_RF 0.147 iacontact local-sim RA 0.608 AA 0.516 JC 0.118 CN 0.505 CN 0.123 global-sim PA 0.592 PA 0.517 PA 0.112 PA 0.505 PA 0.112 learning GS_LSTM_Avg_LR 0.560 GS_LSTM_L1_LR 0.573 GS_MeanPool_Hada_LR 0.103 GS_LSTM_L1_RF 0.595 GS_MaxPool_Avg_RF 0.123 forum local-sim AA 0.617 RA 0.629 CN 0.167 RA 0.590 CN 0.198 global-sim PA 0.732 PA 0.698 PA 0.323 PA 0.658 PA 0.260 learning DW_L2_LR 0.588 DW_Hada_RF 0.580 DW_L2_RF 0.137 DW_L2_LR 0.602 DW_L1_RF 0.167 topo local-sim RA 0.848 RA 0.897 RA 0.603 RA 0.811 RA 0.614 global-sim PA 0.922 PA 0.911 PA 0.608 PA 0.831 PA 0.697 learning GS_MaxPool_Hada_RF 0.818 GS_MaxPool_Hada_RF 0.804 DW_Avg_RF 0.410 GS_MaxPool_Hada_RF 0.743 DW_Avg_RF 0.408 act-mooc local-sim RA 0.840 RA 0.834 RA 0.398 RA 0.771 AA 0.378 global-sim PA 0.994 PA 0.991 PA 0.915 PA 0.984 PA 0.894 learning DW_Avg_LR 0.989 DW_Avg_LR 0.984 DW_Avg_LR 0.840 N2V_L1_LR 0.975 DW_Avg_LR 0.829 arxiv local-sim RA 0.951 RA 0.952 RA 0.662 RA 0.900 RA 0.677 global-sim Norm-CT 0.854 CT 0.828 CT 0.339 CT 0.793 Katz 0.427 learning N2V_L2_LR 0.797 GS_Mean_L1_LR 0.819 DW_L1_LR 0.376 GS_MeanPool_L2_LR 0.745 DW_L1_LR 0.421 pow-grid local-sim CN 0.570 CN 0.772 CN 0.314 AA 0.574 CN 0.179 global-sim Katz 0.606 Katz 0.682 Katz 0.195 PA 0.617 Katz 0.265 learning GS_MaxPool_L1_LR 0.616 GS_MaxPool_L2_LR 0.693 GS_MaxPool_L1_LR 0.157 GS_Mean_L1_LR 0.679 GS_MaxPool_L2_RF 0.240 routers local-sim AA 0.833 AA 0.871 AA 0.431 AA 0.803 AA 0.452 global-sim Katz 0.760 Katz 0.834 Katz 0.438 Katz 0.785 Katz 0.433 learning GS_LSTM_L1_LR 0.690 DW_L2_RF 0.738 DW_L2_LR 0.204 GS_Mean_L2_LR 0.734 DW_L1_RF 0.259 bio-yeast local-sim RA 0.877 RA 0.882 RA 0.525 AA 0.805 RA 0.535 global-sim Katz 0.863 Katz 0.860 Katz 0.465 Katz 0.781 Katz 0.471 learning DW_L2_LR 0.852 DW_L2_LR 0.863 DW_L2_LR 0.466 DW_L2_LR 0.765 DW_L2_LR 0.467 fb local-sim RA 0.945 RA 0.956 RA 0.630 RA 0.925 RA 0.636 global-sim Norm-CT 0.903 Katz 0.919 Katz 0.528 Katz 0.886 Katz 0.543 learning DW_L2_LR 0.914 DW_L1_RF 0.928 DW_L1_LR 0.554 DW_L2_LR 0.888 DW_L2_LR 0.582 blog-cat local-sim RA 0.932 RA 0.935 RA 0.675 RA 0.872 RA 0.656 global-sim PA 0.921 PA 0.925 PA 0.642 PA 0.885 PA 0.625 learning DW_L1_LR 0.949 DW_Avg_LR 0.903 DW_L1_LR 0.723 DW_Avg_LR 0.834 DW_Avg_LR 0.581
DATASET METHOD TYPE AUROC AUPR Pr@P balanced imbalanced balanced imbalanced Method Name Value Method Name Value Method Name Value Method Name Value Method Name Value clg-msg-t global-sim PA 0.753 PA 0.737 PA 0.268 PA 0.685 Katz 0.316 learning GS_MeanPool_Avg_RF 0.696 DW_Avg_LR 0.688 GS_MeanPool_Avg_RF 0.187 GS_MeanPool_Avg_RF 0.648 DW_Avg_RF 0.243 email-t global-sim Katz 0.742 Katz 0.720 Katz 0.218 Katz 0.658 Katz 0.286 learning GS_Mean_L1_RF 0.793 GS_Mean_L1_RF 0.794 GS_MeanPool_Hada_RF 0.326 GS_Mean_L1_LR 0.748 GS_LSTM_L1_LR 0.396 math-flow-t global-sim PA 0.888 PA 0.883 PA 0.531 PA 0.806 PA 0.525 learning N2V_Avg_LR 0.812 DW_Avg_LR 0.811 Dw_Avg_LR 0.381 DW_Avg_LR 0.742 DW_Avg_LR 0.408 ast-phys-t global-sim Katz 0.941 Katz 0.944 Katz 0.754 Katz 0.973 PA 0.684 learning N2V_Avg_RF 0.987 N2V_Avg_RF 0.983 N2V_Avg_RF 0.919 N2V_L2_LR 0.990 N2V_Avg_RF 0.833 lst-fm-t global-sim CT 0.504 PA 0.733 PA 0.195 PA 0.913 Norm-CT 0.168 learning GS_Mean_Avg_RF 0.734 GS_Mean_L2_LR 0.723 GS_MaxPool_Hada_RF 0.264 GS_Mean_Hada_RF 0.940 GS_Mean_L2_LR 0.309 forum-t global-sim PA 0.706 PA 0.672 PA 0.185 PA 0.638 PA 0.227 learning DW_Avg_LR 0.649 DW_L2_LR 0.629 DW_L1_LR 0.153 DW_Avg_LR 0.610 DW_L1_LR 0.175 topo-t global-sim PA 0.905 PA 0.906 PA 0.550 PA 0.837 PA 0.533 learning GS_MeanPool_Avg_RF 0.906 GS_MaxPool_Avg_RF 0.903 GS_MeanPool_Hada_RF 0.609 GS_MaxPool_Avg_RF 0.827 GS_MaxPool_Avg_RF 0.609 act-mooc-t global-sim PA 0.967 PA 0.983 PA 0.910 PA 0.969 PA 0.893 learning N2V_L1_LR 0.964 N2V_L1_LR 0.966 N2V_L1_LR 0.902 N2V_L1_LR 0.969 N2V_L1_LR 0.893
DATASET METHOD TYPE AUROC AUPR Pr@P balanced imbalanced balanced imbalanced Method Name Value Method Name Value Method Name Value Method Name Value Method Name Value clg-msg global-sim PA 0.887 PA 0.887 PA 0.527 PA 0.823 PA 0.535 learning DW_L2_LR 0.648 DW_L1_RF 0.669 DW_L1_RF 0.178 DW_L2_LR 0.640 DW_L1_RF 0.235 email global-sim Katz 0.611 Norm-CT 0.775 Katz 0.151 Katz 0.583 Katz 0.250 learning GS_Attentional_Avg_LR 0.675 GS_MaxPool_L1_LR 0.722 GS_LSTM_L2_RF 0.252 N2V_Hada_RF 0.666 GS_Mean_L1_RF 0.333 math-flow global-sim PA 0.886 PA 0.869 PA 0.489 PA 0.952 PA 0.476 learning GS_MeanPool_Hada_LR 0.719 GS_MeanPool_Hada_LR 0.714 GS_MeanPool_L1_LR 0.224 GS_MaxPool_L2_LR 0.928 GS_MaxPool_L1_LR 0.291 ast-phys global-sim PA 0.991 PA 0.996 Katz 0.906 PA 0.971 PA 0.857 learning N2V_L2_RF 0.924 GS_MaxPool_L1_RF 0.787 N2V_L2_RF 0.584 N2V_L1_LR 0.971 N2V_L2_RF 0.571 forum global-sim Katz 0.742 Katz 0.779 Katz 0.250 Katz 0.682 Katz 0.346 learning DW_L2_LR 0.625 GS_MaxPool_Hada_LR 0.601 DW_L2_LR 0.127 DW_L2_LR 0.567 DW_L1_RF 0.173 topo global-sim PA 0.879 PA 0.873 CT 0.492 PA 0.779 Norm-CT 0.485 learning GS_MeanPool_Hada_RF 0.448 GS_Mean_Hada_RF 0.880 GS_MeanPool_Hada_RF 0.549 GS_MeanPool_Hada_RF 0.805 GS_Mean_Hada_RF 0.552 pow-grid global-sim Norm-CT 0.616 CT 0.594 CT 0.119 Norm-CT 0.602 -Norm-CT 0.136 learning GS_Mean_L2_RF 0.596 GS_Mean_L2_RF 0.723 GS_MaxPool_L1_LR 0.126 GS_Mean_L1_RF 0.630 N2V_L1_RF 0.150 routers global-sim Norm-CT 0.748 Norm-CT 0.809 Norm-CT 0.288 CT 0.716 CT 0.333 learning GS_MaxPool_Hada_LR 0.720 DW_Hada_LR 0.711 GS_MaxPool_L1_LR 0.216 GS_LSTM_L2_LR 0.641 GS_MaxPool_L1_LR 0.320 boi-yeast global-sim PA 0.762 PA 0.715 CT 0.244 PA 0.933 Norm-CT 0.311 learning DW_L2_LR 0.721 DW_L1_LR 0.696 GS_LSTM_Hada_RF 0.231 DW_L2_LR 0.933 GS_Mean_Avg_RF 0.255 fb global-sim Norm-CT 0.942 Katz 0.944 -NormCT 0.700 PA 0.875 Katz 0.250 learning N2V_L1_LR 0.948 GS_MaxPool_L2_RF 0.935 GS_MaxPool_Hada_RF 0.713 DW_Hada_LR 0.100 DW_Hada_LR 0.625 blog-cat global-sim PA 0.807 PA 0.801 Norm-CT 0.392 PA 0.947 PA 0.391 learning N2V_Avg_RF 0.792 N2V_Avg_LR 0.742 N2V_Avg_RF 0.310 N2V_Avg_RF 0.939 N2V_Avg_RF 0.347

### v2

@@
 We identify a number of such factors, such as, network-type, problem-type, geodesic distance between the end nodes and its distribution over the classes, nature and applicability of LP methods, class imbalance and its impact on early retrieval, evaluation metric, etc., and present an experimental setup which allows us to evaluate LP methods in a rigorous and controlled manner.
-We perform extensive experiments with a variety of LP methods over real network datasets in this controlled setup, and gather valuable insights on the interactions of these factors with the performance of LP through an array of carefully designed hypotheses.
-Following the insights, we provide recommendations to be followed as best practice for evaluating LP methods.
+We perform extensive experiments with 58 LP methods of diverse categories over 24 real network datasets in this controlled setup, and gather valuable insights on the interactions of these factors with the performance of LP through an array of carefully designed hypothesis tests.
+Following the insights, we provide recommendations on the experimental setup and the evaluation measures to be employed considering application scenarios, to be followed as best practice for evaluating LP methods.
 keywords:
@@
 We provide the details of the LP methods which we use in our experiments in Section 3, grouped by their categories.
-Section 4 provides the details of the datasets on which we apply the LP methods.
-We describe the proposed experimental setup in Section 5.
+Section 4 provides the details of the datasets on which we apply the LP methods along with the proposed experimental setup.
 In Section 5, we present the results and analyze our findings, and provide our recommendations.
-We conclude our work and indicate the future directions in Section 7.
+We conclude our work and indicate the limitations and future directions in Section 6.
 2 Related Works
@@
 In this paper, we consider the unsupervised methods in our experiments.
-The main reason to exclude the end-to-end methods is the difficulty in building the experimental set-up for those in the longitudinal setup for the future LP, which incurs a longitudinal bias [3, 35] on the prediction performance, as well as makes the methods incomparable with their missing LP counterpart.
+The main reason to exclude the end-to-end methods is the difficulty in building the experimental set-up for those in a longitudinal setup, for the future LP, which incurs a longitudinal bias [3, 35] on the prediction performance, as well as makes the methods incomparable with their missing LP counterpart.
 Another group of unsupervised LP approaches are matrix factorization based [23, 36], which we do not include in our experiments.
@@
 So, excluding a few prediction methods does not hurt the merit of this work.
-The start-of-the-art LP methods briefed above mainly apply to a simple, undirected and homogeneous networks.
-There exists works which exploit several network attributes to predict links.
+The state-of-the-art LP methods briefed above mainly apply to a simple, undirected and homogeneous networks.
+There exist works which exploit several network attributes to predict links.
 Few examples are: temporal link prediction [7, 25] which exploits temporal dynamics of link maintenance in dynamic networks like in social networks; link prediction in bipartite networks [37, 38] which applies on bipartite graphs like user-product networks or term-document networks; link prediction in heterogeneous networks [7, 39, 40] which exploits the heterogeneity of nodes and links in networks like bibliographic networks; link prediction in directed networks [41, 42] which exploits the direction information in directed networks like email networks, etc.
 These specialized methods are not necessarily disjoint, and often use information covering multiple groups.
-In this paper, we consider the start-of-the-art LP methods only in our analysis, and reserve the specialized methods for future study.
+In this paper, we consider the state-of-the-art LP methods only in our analysis, and reserve the specialized methods for future study.
 Studies like [43, 35] investigated the issues associated with LP evaluation in considerable details.
@@
 In total, our experiments are conducted on 58 distinct LP methods, of which, 4 are local-sim, 6 are global-sim, and 48 are learning based methods.
-4 Datasets
-We conducted our experiments on fifteen real network datasets of diverse nature, such as, social network, communication network, collaboration network, biological network, power grid, etc.
+| Network | #Nodes | #Edges | Type | Time Span | Dataset | #Positives | #two-hop positives | #three-hop positives |
+| CollegeMsg [57] | 1899 | 20296 | Directed | 194 days | clg-msg-t | 2142 | 1068(49.80%) | 983(45.89%) |
+| clg-msg | 510 | 335(65.68%) | 174(34.11%) | |||||
+| Email-Eu [58] | 986 | 332334 | Directed | 803 days | email-t | 3977 | 3774(94.89%) | 200(5.03%) |
+| 441 | 429(97.27%) | 11(2.49%) | ||||||
+| Mathoverflow [58] | 24818 | 506550 | Directed | 2350 days | math-flow-t | 24225 | 18918(78.09%) | 4889(20.18%) |
+| math-flow | 11373 | 9981(87.76%) | 1125(9.89%) | |||||
+| Ast-Phys [59] | 9756 | 86697 | Undirected | 3316 days | ast-phys-t | 32186 | 27035(83.99%) | 5151(16%) |
+| ast-phys | 8898 | 8862(99.59%) | 36(0.40%) | |||||
+| LastFm [60] | 7624 | 27806 | Undirected | 1345 days | lst-fm-t | 29867 | 29862(99.38%) | 185(0.61%) |
+| lst-fm | 3755 | 3748(99.81%) | – | |||||
+| IAcontact [61] | 113 | 2498 | Directed | 1 day | iacontact-t | 533 | 533(100%) | – |
+| iacontact | 92 | 92(100%) | – | |||||
+| FB-forum [61] | 899 | 7089 | Directed | 164 days | forum-t | 1638 | 717(43.58%) | 862(52.81%) |
+| forum | 277 | 163(58.84%) | 109(39.35%) | |||||
+| Topology [62, 63] | 34761 | 171403 | Undirected | 23 days | topo-t | 21100 | 17331(82.13%) | 3159(14.97%) |
+| topo | 4682 | 3578(76.42%) | 980(20.93%) | |||||
+| Mooc [60] | 7047 | 411749 | Directed | 29 days | act-mooc-t | 38448 | 38381(99.82%) | 67(0.01%) |
+| act-mooc | 8594 | 8590(99.95%) | – | |||||
+| ArXiv [59] | 23133 | 93133 | Undirected | arxiv | 886 | 867(97.85%) | 6(0.14%) | |
+| Power-grid [61] | 4941 | 6594 | Undirected | pow-grid | 183 | 117(63.93%) | 65(35.51%) | |
+| Tech-routers [61] | 2113 | 6632 | Undirected | routers | 590 | 72(12.20%) | 45(7.62%) | |
+| Bio-yeast [61] | 6008 | 156945 | Undirected | bio-yeast | 1572 | 1472(93.63%) | 100(6.36%) | |
+| Facebook [64] | 4039 | 88234 | Undirected | fb | 8813 | 8796(99.80%) | 17(0.19%) | |
+| BlogCatalog [61] | 10312 | 333983 | Undirected | blog-cat | 33350 | 33112(99.28%) | 238(0.71%) |
+4 Experimental Setup
+4.1 Network datasets
+We conducted our experiments on fifteen real network datasets of diverse nature, such as social network, communication network, collaboration network, biological network, power grid, etc.
 A few of these networks like friendship networks, collaboration networks are undirected in nature, whereas networks like communication networks are directed.
@@
 In total, nine among all the networks had timestamps.
-For these networks we created ground truth labels two ways: (a) considering the timestamps which lets us evaluate future LP, and (b) removing a portion of the links in the networks which lets us evaluate the missing LP.
+For these networks we created ground truth labels in two ways: (a) considering the timestamps which lets us evaluate future LP, and (b) removing a portion of the links in the networks which lets us evaluate the missing LP.
 So, we created two LP datasets for each of these networks.
@@
 We summarize the datasets in Table 2.
-We briefly describe the datasets as follows.
-4.1 Networks with timestamps:
-1.
-CollegeMsg [57]: This dataset represents a temporal network capturing private communications among college students on a campus at the University of California, Irvine.
-Each node in the network represents an individual student, while the private messages exchanged between them guides the link formation and maintenance.
-This dataset offers insights into communication dynamics and prevalent styles within the campus community.
-Information for this dataset was obtained over the course of 193 days at the University of California, Irvine.
-In this paper, we refer the dataset which prepares ground truth using timestamps as clg-msg-t and the the one which do not consider the timestamps as clg-msg.
-2.
-Email-Eu [58]: This dataset provides insight into the evolving connections and communication patterns within a European research organization over time.
-It primarily focuses on email correspondence exchanged between members of the institution.
-The network’s nodes represent individual users or email addresses within the organization, and the links encode the email exchanges between them, with timestamps indicating the exact moments of interaction.
-This data enables researchers to analyze the dynamics of information flow and collaboration within the organization.
-Information for this dataset was gathered over a period of 803 days.
-In this paper, we refer the temporal version as email-t and the non-temporal one as email.
-3.
-Mathoverflow [58]: This dataset presents a temporal network capturing interactions among users on the Mathoverflow platform, a specialized community for mathematicians to exchange knowledge and insights.
-Each node in the network represents an individual user, while the links encode their interactions, such as comments or responses to questions and answers.
-The timestamps associated with the interactions indicate the exact moment of these interactions, providing valuable temporal insights into the dynamics of knowledge sharing and collaboration within the mathematical research community.
-This dataset is taken over a time span of 2350 days.
-In this paper, we refer the temporal version as math-flow-t and the non-temporal version as math-flow.
-4.
-Ast-phys [59]: This dataset contains a collaboration network of authors who have contributed to the field of Astrophysics through scientific paper publications.
-Each node represents an individual author, and the links signify collaborations between authors.
-The timestamps indicate the time of collaborations between two authors.
-In this paper, we refer the temporal version as ast-phys-t and the non-temporal version as ast-phys.
-5.
-LastFm [60]: This dataset represents a social network derived from Last.fm, a music streaming and recommendation service.
-Nodes in the network represent Last.fm users, and the links symbolize their interactions and social relationships.
-The dataset includes timestamps indicating the precise moments of user connections and interactions.
-In this paper, we refer the temporal version as lst-fm-t and the non-tempral one as lst-fm.
-6.
-IAcontact [61]: This dataset represents a contact network among attendees of the ACM Hypertext conference in the year 2009.
-The links encode the interactions among the attendees during a 2.5-day period, while the nodes represent the attendees.
-In this paper, we refer the temporal version as iacontact-t and the non-temporal one as iacontact.
-7.
-FB-forum [61]: This dataset illustrates a forum-based online social network similar to Facebook.
-The network mostly records user actions and conversations in this forum.
-Individual nodes in this dataset represent users, and the links represent the forum messages these users have exchanged.
-The dataset spans a period of 164 days.
-In this paper, we refer the temporal version as forum-t and the non-temporal one as forum.
-8.
-Topology [62, 63]: This dataset encodes the complicated network of connections between autonomous systems on the Internet.
-Node refers to autonomous systems (AS), which are collections of connected IP routing domains run by independent network operators.
-In this paper, we refer the temporal version as topo-t and the non-temporal one as topo.
-9.
-Mooc [60]: This dataset displays user activity on a well-known MOOC program.
-The actions are symbolized as a temporal and directed network.
-Users are represented as nodes in the graph together with course activities (targets), and the links are user actions on the targets.
-In this paper, we refer the temporal version as act-mooc-t and the non-temporal one as act-mooc.
-4.2 Networks without timestamps:
-1.
-ArXiv [59]: ArXiv is a network dataset representing collaborations among scientists who publish preprints and papers on the ArXiv platform.
-Nodes in the network represent individual authors, while links represent co-authorship connections.
-We refer the dataset prepared from this network as arxiv.
-2.
-Power-grid [61]: The Power-grid network dataset models the interconnected power grid infrastructure in the United States.
-It provides information on the transmission lines, nodes, and elements that make up the electrical system, enabling research into electricity distribution across different geographic areas.
-Nodes represent power plants, substations, transformers, and other infrastructure, while links depict the transmission lines connecting these elements.
-This dataset is valuable for studying the dynamics and efficiency of the U.S.
-electrical power grid.
-In this paper, we refer the dataset prepared from this network as pow-grid.
-3.
-Tech-routers [61]: It is a tech routers network, which represents the connections between routers in a technological infrastructure like the internet or computer networks.
-Nodes in the network represent individual routers, and links symbolize direct physical or logical links between matching routers.
-We refer the dataset prepared from this network as routers.
-4.
-Bio-yeast [61]: The Bio-yeast dataset is a biological network database focused on protein-protein interactions in the yeast species Saccharomyces cerevisiae.
-It represents proteins as nodes and shows connections between proteins based on physical binding, enzymatic processes, or co-expression patterns.
-We refer the dataset prepared from this network as bio-yeast.
-5.
-Facebook [64]: It is a social network, built on the Facebook platform, focuses on a single user (“ego”) and their connections with other users.
-Nodes represent Facebook users, and links represent friendships between them.
-We refer the dataset prepared from this network as fb.
-6.
-BlogCatalog [61]: This dataset is a social relationship network.
-The network is made up of bloggers and their friends and other social connections.
-Nodes in this network are represented as bloggers and the links represent social relationship among the bloggers.
-In this paper, we refer the dataset prepared from this network as blog-cat.
-| Network | #Nodes | #Edges | Type | Time Span | Dataset | #Positives | #two-hop positives | #three-hop positives |
-| CollegeMsg | 1899 | 20296 | Directed | 194 days | clg-msg-t | 2142 | 1068(49.80%) | 983(45.89%) |
-| clg-msg | 510 | 335(65.68%) | 174(34.11%) | |||||
-| Email-Eu | 986 | 332334 | Directed | 803 days | email-t | 3977 | 3774(94.89%) | 200(5.03%) |
-| 441 | 429(97.27%) | 11(2.49%) | ||||||
-| Mathoverflow | 24818 | 506550 | Directed | 2350 days | math-flow-t | 24225 | 18918(78.09%) | 4889(20.18%) |
-| math-flow | 11373 | 9981(87.76%) | 1125(9.89%) | |||||
-| Ast-Phys | 9756 | 86697 | Undirected | 3316 days | ast-phys-t | 32186 | 27035(83.99%) | 5151(16%) |
-| ast-phys | 8898 | 8862(99.59%) | 36(0.40%) | |||||
-| LastFm | 7624 | 27806 | Undirected | 1345 days | lst-fm-t | 29867 | 29862(99.38%) | 185(0.61%) |
-| lst-fm | 3755 | 3748(99.81%) | – | |||||
-| IAcontact | 113 | 2498 | Directed | 1 day | iacontact-t | 533 | 533(100%) | – |
-| iacontact | 92 | 92(100%) | – | |||||
-| FB-forum | 899 | 7089 | Directed | 164 days | forum-t | 1638 | 717(43.58%) | 862(52.81%) |
-| forum | 277 | 163(58.84%) | 109(39.35%) | |||||
-| Topology | 34761 | 171403 | Undirected | 23 days | topo-t | 21100 | 17331(82.13%) | 3159(14.97%) |
-| topo | 4682 | 3578(76.42%) | 980(20.93%) | |||||
-| Mooc | 7047 | 411749 | Directed | 29 days | act-mooc-t | 38448 | 38381(99.82%) | 67(0.01%) |
-| act-mooc | 8594 | 8590(99.95%) | – | |||||
-| ArXiv | 23133 | 93133 | Undirected | arxiv | 886 | 867(97.85%) | 6(0.14%) | |
-| Power-grid | 4941 | 6594 | Undirected | pow-grid | 183 | 117(63.93%) | 65(35.51%) | |
-| Tech-routers | 2113 | 6632 | Undirected | routers | 590 | 72(12.20%) | 45(7.62%) | |
-| Bio-yeast | 6008 | 156945 | Undirected | bio-yeast | 1572 | 1472(93.63%) | 100(6.36%) | |
-| 4039 | 88234 | Undirected | fb | 8813 | 8796(99.80%) | 17(0.19%) | ||
-| BlogCatalog | 10312 | 333983 | Undirected | blog-cat | 33350 | 33112(99.28%) | 238(0.71%) |
-5 Experimental Setup
+Note that, the column titled “Network” gives the network names, where the “Dataset” column gives the names of the datasets for link prediction after generating the graph and ground truth labels.
+The first nine networks had timestamps, so each produces two datasets.
+The name of the datasets which consider timestamps, are suffixed with -t.
+For an example, the “CollegeMsg” network produces two datasets; the timestamped one is named as clg-msg-t, and the non-timestamped one is named as clg-msg.
+We present a brief description of the networks at Appendix A.
 4.2 Graph and ground truth generation
-In this section, we detail the procedure of how we prepare the graph and the ground truth positive and negative node pairs from the datasets presented in Section 4, for future and missing LP tasks.
+In this section, we detail the procedure of how we prepare the graph and the ground truth positive and negative node pairs from the network datasets presented in Section 4.1, for future and missing LP tasks.
 Unlike traditional approaches, we carefully segregate the test examples based on the distance of the test node pairs in the graph, and maintain the distance bias inherent with the positive examples’ distribution, while sampling the negative examples.
@@
 From this potential set, we first ignored the node pairs which involve at least one node which is not included in the graph, and then removed the links which are present in the graph as an edge to get the final set of positive examples.
-The missing LP was applied to all the networks presented in Section 4.
+A pseudocode for the graph and positive examples generation for future LP is given in Algorithm 1.
+training_interactions first 75%
+future_interactions last 25%
+The missing LP was applied to all the networks presented in Section 4.1.
 We preprocessed the networks having timestamp information by ignoring the timestamps and their directions, and then removing the duplicates.
@@
 We consider the removed links as the potential positive examples, and generate the positive examples following the same procedure as the future LP datasets.
+A pseudocode for the graph and positive examples generation for missing LP is given in Algorithm 2.
 For each dataset, we divided the positive examples in three parts based on the shortest path distance of the end node pair in the graph: having the shortest path distance (a) , (b) , and (c) more than .
@@
 Despite AUPR’s suitability, it is less popular in LP literature, may be due to the variability of PR curves in different skew levels, which makes it difficult to compare the LP methods when multiple skewness is considered.
-To deal with class imbalance, another skew sensitive evaluation measure, is sometimes used to evaluate LP as a single threshold measure, where the positive example count is considered as a typical value of [65, 2].
+Measures like , , F-score, etc.
+are less popular in link prediction literature, mainly because, unlike AUROC and AUPR, they measure performance for a single threshold value, which prohibits us to understand the method’s behavior throughout the whole spectrum of thresholds [3, 2, 24].
+, another skew sensitive evaluation measure, is sometimes used to evaluate LP as a single threshold measure, where the positive example count is considered as a typical value of [65, 2].
 In this work, we consider AUROC, AUPR and as single point summary metrics for evaluating LP methods.
@@
 5 Results and Discussion
-We applied the 58 LP methods presented in Section 3 on the 24 datasets defined in Section 4 in a controlled experimental setup proposed in Section 5.
-We evaluated their performance based on the four evaluation metrics: AUROC, AUPR, and on the test sets.
-As described in Section 5.1, we had different test sets depending on the shortest path distance between test node pairs, and each of these had two versions: balanced and imbalanced.
-For the skew sensitive metrics: AUPR, and , we consider evaluating both for balanced and imbalanced test set.
+We applied the 58 LP methods presented in Section 3 on the 24 datasets defined in Section 4.1 in a controlled experimental setup proposed in Section 4.
+We evaluated their performance based on the evaluation metrics presented in Section 4.4.
 We performed a series of statistical tests to test an array of carefully designed hypotheses in this controlled setup to interrogate how the LP methods perform in different situations depending on the shortest path distance between test node pairs, prediction type, network type, and class imbalance.
 We also investigate how the class imbalance affect the LP methods in early retrieval phase using proper statistical tests and PRC plots.
-We present the detailed results along with the insights acquired from them as follows.
+We present the detailed results along with the insights acquired from them, and our recommendations as follows.
 5.1 Future LP Vs. missing LP
@@
 Given a network, we apply paired t-test to compare the absolute values of the prediction performances of the prediction methods given a particular evaluation metric for the two cases.
-To find the level of the ordinal association between the two cases, we apply Kendall?s Tau Rank Correlation Coefficient ().
+To find the level of the ordinal association between the two cases, we apply Kendall’s Tau Rank Correlation Coefficient ().
 We consider AUROC, and both balanced and imbalanced versions of AUPR and as the evaluation measure.
 We perform the experiment on the nine networks which had timestamps in distance controlled test set.
-Table 3 and Table 4 summarize the results of the two tests respectively.
+Table 3 summarizes the paired t-test results.
 Almost all paired t-test p-values are significant across distances and evaluation measures, and for almost all networks, the LP methods perform better for future LP than missing LP.
@@
 This also indicates that when an application demands future LP, but the ground truth links are generated in traditional way, i.e., removing a portion of links from the network randomly, the prediction performance is significantly underestimated, and vice versa.
-Table 4 shows that except for the Topology and Mooc network, all rank correlation values are extremely low, in few cases even negative.
+Kendall’s Tau results reveal that except for the Topology and Mooc network, all rank correlation values are extremely low, in few cases even negative.
 This indicates that, for a future LP, if the ground truth links are generated in traditional way, the prediction performance rankings of various LP methods vary significantly than if the ground truth links were generated with true future links.
+All correlation values can be found in Table 16 in Appendix C.
 Our results indicate the necessity to evaluate a LP method in both setups and report its performance in both types of prediction tasks, in order to understand its effectiveness.
@@
 | Mooc | — | — | — | — | — | — | — | — | — | — | |
-| NETWORK | AUROC | AUPR | Pr@P | |||
-| balanced | imbalanced | balanced | imbalanced | |||
-| two-hop | CollegeMsg | 0.184 | 0.184 | 0.219 | 0.156 | 0.203 |
-| Email-Eu | 0.272 | 0.319 | 0.472 | 0.294 | 0.526 | |
-| Mathoverflow | 0.339 | 0.386 | 0.414 | 0.326 | 0.488 | |
-| Ast-phys | 0.487 | 0.433 | 0.377 | 0.525 | 0.393 | |
-| lastFm | 0.477 | 0.465 | 0.381 | 0.462 | 0.335 | |
-| IAcontact | 0.027 | 0.080 | 0.038 | 0.039 | 0.075 | |
-| FB-forum | 0.321 | 0.347 | 0.438 | 0.311 | 0.271 | |
-| Topology | 0.785 | 0.783 | 0.771 | 0.723 | 0.742 | |
-| Mooc | 0.744 | 0.743 | 0.698 | 0.678 | 0.741 | |
-| three-hop | CollegeMsg | 0.009 | 0.118 | 0.158 | 0.022 | 0.230 |
-| Email-Eu | -0.030 | 0.085 | 0.167 | -0.144 | 0.181 | |
-| Mathoverflow | -0.020 | -0.071 | 0.018 | 0.139 | 0.060 | |
-| Ast-phys | 0.256 | 0.464 | 0.578 | 0.422 | 0.628 | |
-| lastFm | — | — | — | — | — | |
-| IAcontact | — | — | — | — | — | |
-| FB-forum | 0.189 | 0.208 | 0.283 | 0.203 | 0.334 | |
-| Topology | 0.749 | 0.747 | 0.738 | 0.649 | 0.773 | |
-| Mooc | — | — | — | — | — | |
 5.2 Distance between the end nodes in a test node pair
@@
 For each dataset, we performed paired t-test to determine if the prediction performance differ in terms of their absolute values for the two cases: two-hop away test node pairs and three-hop away test node pairs.
-Also, we measure the Kendall?s values for the the same to estimate the ordinal association between the two cases.
+Also, we measure the Kendall’s values for the same to estimate the ordinal association between the two cases.
 We consider AUROC, AUPR (balanced), AUPR (imbalanced), (balanced) and (imbalanced) as the evaluation measures.
 We performed the experiment across all datasets and prediction types.
-We summarize the results of the two tests in Table 5 and Table 6 respectively.
 Table 4 shows that the absolute performance of the prediction methods differs significantly in most of the cases across all datasets, prediction types and evaluation metrics.
@@
 Moreover, this behavior differs over the prediction types.
-From Table 6 it is also evident that the rankings of the prediction methods for the two cases vary significantly.
+From the Kendall’s Tau results it is also evident that the rankings of the prediction methods for the two cases vary significantly.
+All correlation values can be found in Table 17 at Appendix C.
 Overall, the obtained results support the need to evaluate LP methods separately for the two-hop and three-hop case to understand its behavior in distance controlled manner, and our experimental setup provides an opportunity to measure it.
-Moreover, as the local LP methods are applicable only on in the two-hop case, but the global methods are applicable in both the cases, our experimental setup provides a robust environment to compare a local with a global method.
+Moreover, as the local LP methods are applicable only on the two-hop case, but the global methods are applicable in both the cases, our experimental setup provides a robust environment to compare a local with a global method.
 | DATASET | AUROC | AUPR | Pr@P | ||||||||
@@
 | act-mooc-t | 0.265 | — | .004** | 2-hop | 0.824 | — | .001*** | 3-hop | 0.372 | — | |
-| DATASET | AUROC | AUPR | Pr@P | |||
-| balanced | imbalanced | balanced | imbalanced | |||
-| missing LP | clg-msg | 0.275 | 0.203 | 0.241 | 0.316 | 0.207 |
-| -0.059 | -0.012 | -0.155 | -0.091 | 0.007 | ||
-| math-flow | 0.399 | 0.385 | 0.227 | 0.311 | 0.196 | |
-| ast-phys | 0.278 | 0.391 | 0.572 | 0.373 | 0.542 | |
-| lst-fm | — | — | — | — | — | |
-| forum | 0.184 | 0.227 | 0.236 | 0.221 | 0.221 | |
-| topo | 0.233 | 0.316 | 0.302 | 0.176 | 0.241 | |
-| act-mooc | — | — | — | — | — | |
-| pow-grid | -0.144 | 0.084 | 0.173 | -0.108 | 0.189 | |
-| routers | 0.242 | 0.005 | 0.079 | 0.285 | 0.068 | |
-| bio-yeast | 0.216 | 0.260 | 0.362 | 0.316 | 0.302 | |
-| fb | -0.178 | 0.450 | 0.211 | -0.159 | 0.025 | |
-| blog-cat | 0.213 | 0.031 | 0.049 | 0.319 | 0.033 | |
-| future LP | clg-msg-t | 0.580 | 0.589 | 0.548 | 0.578 | 0.489 |
-| email-t | 0.603 | 0.443 | 0.390 | 0.609 | 0.337 | |
-| math-flow-t | 0.387 | 0.437 | 0.467 | 0.391 | 0.445 | |
-| ast-phys-t | 0.660 | 0.705 | 0.677 | 0.508 | 0.653 | |
-| lst-fm-t | 0.112 | -0.021 | -0.095 | 0.222 | -0.131 | |
-| forum-t | 0.476 | 0.467 | 0.511 | 0.488 | 0.406 | |
-| topo-t | 0.363 | 0.391 | 0.360 | 0.326 | 0.361 | |
-| act-mooc-t | 0.439 | 0.589 | 0.573 | 0.413 | 0.594 | |
 5.3 Prediction method type
@@
 In Table 9, we report the test results.
-For the best performing methods in each group given an evaluation metric, readers may refer to Appendix A.
+For the best performing methods in each group given an evaluation metric, readers may refer to Appendix B.
 We find few tests achieve the significance level, and for all such cases the undirected networks are the winner.
@@
 We further do the same with the single threshold imbalance sensitive metric .
-We consider Kendall?s Tau Rank Correlation Coefficient () to measure the (in)consistently of ranks of the LP methods when evaluated in each pair of setups stated above.
+We consider Kendall’s Tau Rank Correlation Coefficient () to measure the (in)consistency of ranks of the LP methods when evaluated in each pair of setups stated above.
 The values for each of the datasets are shown in Table 10.
@@
 We first investigate how the performance on the single point metric (both balanced and imbalanced versions) correlate with that of AUROC, and the balanced and imbalanced versions of correlate with each other, and then show an example ROC plot demonstrating the different behavior of the prediction methods in the two test settings.
-For each dataset, we measure the Kendall?s Tau Rank Correlation Coefficient () of the LP methods for the pair of cases stated above in our controlled setup.
+For each dataset, we measure the Kendall’s Tau Rank Correlation Coefficient () of the LP methods for the pair of cases stated above in our controlled setup.
 Table 11 shows that the correlations with AUROC are significantly decreasing from -balanced to -imbalanced case, which implies that with increasing imbalance, AUROC becomes very inconsistent with the early retrieval behavior.
@@
 DATASET two-hop three-hop AUROC vs Pr@P/2 (balanced) AUROC vs Pr@P/2 (imbalanced) Pr@P/2 (balanced) vs Pr@P/2 (imbalanced) AUROC vs Pr@P/2 (balanced) AUROC vs Pr@P/2 (imbalanced) Pr@P/2 (balanced) vs Pr@P/2 (imbalanced) missing LP clg-msg 0.769 0.514 0.550 0.618 0.369 0.413 email 0.769 0.460 0.380 0.713 0.107 0.185 math-flow 0.836 0.691 0.737 0.874 0.727 0.791 ast-phys 0.840 0.786 0.890 0.571 0.121 0.048 lst-fm 0.736 0.598 0.677 — — — iacontact 0.619 0.212 0.140 — — — forum 0.749 0.566 0.449 0.627 0.477 0.449 topo 0.799 0.776 0.862 0.533 0.455 0.784 act-mooc 0.826 0.761 0.898 — — — arxiv 0.685 0.552 0.692 — — — pow-grid 0.563 -0.009 -0.080 0.622 0.335 0.358 routers 0.812 0.133 0.131 0.717 0.243 0.197 bio-yeast 0.804 0.740 0.830 0.717 0.302 0.387 fb 0.822 0.687 0.721 0.271 0.163 0.181 blog-cat 0.857 0.775 0.877 0.763 0.525 0.503 future LP clg-msg-t 0.815 0.601 0.691 0.817 0.669 0.669 email-t 0.701 0.453 0.604 0.751 0.521 0.601 math-flow-t 0.921 0.801 0.862 0.867 0.773 0.845 ast-phys-t 0.802 0.659 0.824 0.564 0.598 0.703 lst-fm-t 0.798 0.625 0.677 0.642 0.312 0.319 iacontact-t 0.725 0.501 0.461 — — — forum-t 0.785 0.576 0.568 0.830 0.685 0.708 topo-t 0.866 0.795 0.892 0.738 0.623 0.763 act-mooc-t 0.839 0.788 0.913 0.280 0.302 0.161
-6.6 Summary
-Our proposed experimental setup allowed us to conduct the tests presented in this section in order to evaluate the existing LP methods in a controlled environment.
-Our main findings are: (a) LP methods perform better when future LP is considered than that of missing LP; (b) behavior of the LP methods are not consistent on the two test sets: two-hop and three-hop away node pairs; (c) when absolute performance of the LP methods are compared based on their category, we found significant difference in performance for the test set consisting of two-hop away node pairs, but for its three-hop away counterpart, no significant difference was observed; (d) converting directed networks in undirected ones and applying LP methods on it may significantly affect its performance for future LP on the test set of two-hop away node pairs; (e) injecting imbalance in the test set affects the relative performance of the LP methods for the skew sensitive metrics like AUPR and , and the rankings based on AUPR and are not consistent with that of the skew insensitive metric AUROC; (f) rankings of the LP methods differ based on their performance in the early retrieval phase when imbalanced test set used as opposed to balanced one.
-Based on our findings, we recommend evaluating the LP methods in the controlled experimental setup presented in this paper.
-The methods should be evaluated on both the problem types, on the carefully designed distance controlled and skew controlled test set.
-A method’s performance needs to be compared within its category in this setup, as well as with all the methods irrespective of their categories.
-We recommend AUPR as a single point summary evaluation measure to be employed on both of the balanced and imbalanced test set.
-Along with it, PRC, on both of balanced and imbalanced test set, should be investigated to focus on certain portion of the whole spectrum of the thresholds, such as the early retrieval region, which may be necessary for certain applications.
-7 Conclusion and Future Directions
+5.6 Summary of the observation and recommendations
+Our main findings from the experiments are summarized as follows:
+Future LP vs. Missing LP: LP methods perform better when future LP is considered than that of missing LP
+Distance between the end nodes: behavior of the LP methods are not consistent on the two test sets: two-hop and three-hop away node pairs
+Prediction method type: when absolute performance of the LP methods are compared based on their category, we found significant difference in performance for the test set consisting of two-hop away node pairs; but for its three-hop away counterpart, no significant difference was observed
+Network type: converting directed networks to undirected ones and applying LP methods on it may significantly affect its performance of LP, specifically for future LP on the test set of two-hop away node pairs
+Class imbalance: injecting imbalance in the test set affects the relative performance of the LP methods for the skew sensitive metrics like AUPR and , and the rankings based on AUPR and are not consistent with that of the skew insensitive metric AUROC
+Early retrieval: rankings of the LP methods differ based on their performance in the early retrieval phase when imbalanced test set used as opposed to balanced one.
+Based on our findings, we present our recommendations to evaluate a LP method as following.
+6 Conclusion, Limitations, and Future Directions
 In this paper, we identified a number of gaps in the existing link prediction literature concerning its performance evaluation.
@@
 This work can be extended to identify the unexplored avenues in evaluation of LP on specialized graphs, such as weighted, directed and heterogeneous.
-Various aspects of the performance of temporal LP methods and LP methods, which are specifically designed for the directed networks, may be investigated as further studies.
+Various aspects of the performance of temporal LP methods and LP methods that are specifically designed for the directed networks, which this paper did not consider, may be investigated as further studies.
 7 Declarations
@@
 This declaration is “not applicable”.
-Appendix A Consolidated results: Best performing methods in individual groups of methods
+Appendix A Description of Network Datasets
+A.1 Networks with timestamps:
+1.
+CollegeMsg [57]: This dataset represents a temporal network capturing private communications among college students on a campus at the University of California, Irvine.
+Each node in the network represents an individual student, while the private messages exchanged between them guides the link formation and maintenance.
+This dataset offers insights into communication dynamics and prevalent styles within the campus community.
+Information for this dataset was obtained over the course of 193 days at the University of California, Irvine.
+In this paper, we refer the dataset which prepares ground truth using timestamps as clg-msg-t and the the one which do not consider the timestamps as clg-msg.
+2.
+Email-Eu [58]: This dataset provides insight into the evolving connections and communication patterns within a European research organization over time.
+It primarily focuses on email correspondence exchanged between members of the institution.
+The network’s nodes represent individual users or email addresses within the organization, and the links encode the email exchanges between them, with timestamps indicating the exact moments of interaction.
+This data enables researchers to analyze the dynamics of information flow and collaboration within the organization.
+Information for this dataset was gathered over a period of 803 days.
+In this paper, we refer the temporal version as email-t and the non-temporal one as email.
+3.
+Mathoverflow [58]: This dataset presents a temporal network capturing interactions among users on the Mathoverflow platform, a specialized community for mathematicians to exchange knowledge and insights.
+Each node in the network represents an individual user, while the links encode their interactions, such as comments or responses to questions and answers.
+The timestamps associated with the interactions indicate the exact moment of these interactions, providing valuable temporal insights into the dynamics of knowledge sharing and collaboration within the mathematical research community.
+This dataset is taken over a time span of 2350 days.
+In this paper, we refer the temporal version as math-flow-t and the non-temporal version as math-flow.
+4.
+Ast-phys [59]: This dataset contains a collaboration network of authors who have contributed to the field of Astrophysics through scientific paper publications.
+Each node represents an individual author, and the links signify collaborations between authors.
+The timestamps indicate the time of collaborations between two authors.
+In this paper, we refer the temporal version as ast-phys-t and the non-temporal version as ast-phys.
+5.
+LastFm [60]: This dataset represents a social network derived from Last.fm, a music streaming and recommendation service.
+Nodes in the network represent Last.fm users, and the links symbolize their interactions and social relationships.
+The dataset includes timestamps indicating the precise moments of user connections and interactions.
+In this paper, we refer the temporal version as lst-fm-t and the non-tempral one as lst-fm.
+6.
+IAcontact [61]: This dataset represents a contact network among attendees of the ACM Hypertext conference in the year 2009.
+The links encode the interactions among the attendees during a 2.5-day period, while the nodes represent the attendees.
+In this paper, we refer the temporal version as iacontact-t and the non-temporal one as iacontact.
+7.
+FB-forum [61]: This dataset illustrates a forum-based online social network similar to Facebook.
+The network mostly records user actions and conversations in this forum.
+Individual nodes in this dataset represent users, and the links represent the forum messages these users have exchanged.
+The dataset spans a period of 164 days.
+In this paper, we refer the temporal version as forum-t and the non-temporal one as forum.
+8.
+Topology [62, 63]: This dataset encodes the complicated network of connections between autonomous systems on the Internet.
+Node refers to autonomous systems (AS), which are collections of connected IP routing domains run by independent network operators.
+In this paper, we refer the temporal version as topo-t and the non-temporal one as topo.
+9.
+Mooc [60]: This dataset displays user activity on a well-known MOOC program.
+The actions are symbolized as a temporal and directed network.
+Users are represented as nodes in the graph together with course activities (targets), and the links are user actions on the targets.
+In this paper, we refer the temporal version as act-mooc-t and the non-temporal one as act-mooc.
+A.2 Networks without timestamps:
+1.
+ArXiv [59]: ArXiv is a network dataset representing collaborations among scientists who publish preprints and papers on the ArXiv platform.
+Nodes in the network represent individual authors, while links represent co-authorship connections.
+We refer the dataset prepared from this network as arxiv.
+2.
+Power-grid [61]: The Power-grid network dataset models the interconnected power grid infrastructure in the United States.
+It provides information on the transmission lines, nodes, and elements that make up the electrical system, enabling research into electricity distribution across different geographic areas.
+Nodes represent power plants, substations, transformers, and other infrastructure, while links depict the transmission lines connecting these elements.
+This dataset is valuable for studying the dynamics and efficiency of the U.S.
+electrical power grid.
+In this paper, we refer the dataset prepared from this network as pow-grid.
+3.
+Tech-routers [61]: It is a tech routers network, which represents the connections between routers in a technological infrastructure like the internet or computer networks.
+Nodes in the network represent individual routers, and links symbolize direct physical or logical links between matching routers.
+We refer the dataset prepared from this network as routers.
+4.
+Bio-yeast [61]: The Bio-yeast dataset is a biological network database focused on protein-protein interactions in the yeast species Saccharomyces cerevisiae.
+It represents proteins as nodes and shows connections between proteins based on physical binding, enzymatic processes, or co-expression patterns.
+We refer the dataset prepared from this network as bio-yeast.
+5.
+Facebook [64]: It is a social network, built on the Facebook platform, focuses on a single user (“ego”) and their connections with other users.
+Nodes represent Facebook users, and links represent friendships between them.
+We refer the dataset prepared from this network as fb.
+6.
+BlogCatalog [61]: This dataset is a social relationship network.
+The network is made up of bloggers and their friends and other social connections.
+Nodes in this network are represented as bloggers and the links represent social relationship among the bloggers.
+In this paper, we refer the dataset prepared from this network as blog-cat.
+Appendix B Consolidated Results: Best Performing Methods
 DATASET METHOD TYPE AUROC AUPR Pr@P balanced imbalanced balanced imbalanced Method Name Value Method Name Value Method Name Value Method Name Value Method Name Value clg-msg-t local-sim RA 0.678 RA 0.685 RA 0.206 RA 0.630 RA 0.245 global-sim PA 0.751 PA 0.751 PA 0.267 PA 0.693 PA 0.300 learning N2V_Avg_LR 0.688 N2V_Avg_LR 0.688 DW_L2_LR 0.189 N2V_Avg_LR 0.643 N2V_Avg_LR 0.236 email-t local-sim RA 0.860 RA 0.846 RA 0.432 RA 0.786 RA 0.464 global-sim Katz 0.777 Katz 0.753 Katz 0.270 Katz 0.716 Katz 0.310 learning GS_MaxPool_Hada_RF 0.803 DW_L2_LR 0.764 DW_L2_LR 0.278 GS_MaxPool_Hada_RF 0.742 DW_L1_LR 0.348 math-flow-t local-sim RA 0.889 RA 0.902 RA 0.624 RA 0.811 RA 0.589 global-sim PA 0.912 PA 0.911 PA 0.621 PA 0.829 PA 0.579 learning DW_Avg_LR 0.871 DW_Avg_LR 0.860 DW_Avg_LR 0.438 DW_Avg_LR 0.793 DW_Avg_LR 0.465 ast-phys-t local-sim CN 0.789 RA 0.716 RA 0.196 AA 0.721 RA 0.184 global-sim PA 0.941 PA 0.908 PA 0.485 PA 0.856 PA 0.580 learning N2V_Avg_LR 0.634 N2V_Avg_RF 0.974 N2V_Avg_LR 0.756 N2V_Avg_RF 0.945 N2V_Avg_RF 0.736 lst-fm-t local-sim RA 0.591 RA 0.578 RA 0.126 RA 0.565 RA 0.151 global-sim PA 0.598 PA 0.581 PA 0.127 PA 0.570 PA 0.149 learning N2V_Avg_LR 0.576 N2V_Avg_LR 0.565 GS_Mean_Avg_RF 0.120 N2V_Avg_LR 0.557 GS_Mean_Avg_LR 0.148 iacontact-t local-sim RA 0.608 RA 0.608 RA 0.153 AA 0.573 RA 0.180 global-sim PA 0.583 PA 0.584 PA 0.133 PA 0.575 PA 0.167 learning GS_LSTM_L1_RF 0.612 GS_LSTM_L1_RF 0.614 GS_LSTM_L1_RF 0.134 GS_LSTM_L1_RF 0.618 GS_Mean_L2_RF 0.146 forum-t local-sim AA 0.608 AA 0.606 AA 0.146 AA 0.565 AA 0.179 global-sim PA 0.710 PA 0.673 PA 0.193 PA 0.642 PA 0.241 learning DW_Avg_LR 0.641 DW_Avg_RF 0.619 GS_MeanPool_Avg_RF 0.151 DW_Avg_LR 0.596 DW_Avg_RF 0.197 topo-t local-sim AA 0.938 AA 0.945 AA 0.609 AA 0.857 AA 0.581 global-sim PA 0.952 PA 0.951 PA 0.617 PA 0.890 PA 0.588 learning DW_L2_LR 0.914 DW_Avg_LR 0.899 GS_MaxPool_Hada_RF 0.552 DW_Avg_LR 0.832 DW_L1_LR 0.570 act-mooc-t local-sim RA 0.738 RA 0.713 RA 0.23 CN 0.660 RA 0.237 global-sim PA 0.974 PA 0.963 PA 0.763 PA 0.945 PA 0.836 learning DW_Avg_LR 0.977 N2V_Hada_RF 0.976 N2V_Hada_LR 0.857 N2V_Avg_RF 0.945 N2V_L1_LR 0.833
@@
 DATASET METHOD TYPE AUROC AUPR Pr@P balanced imbalanced balanced imbalanced Method Name Value Method Name Value Method Name Value Method Name Value Method Name Value clg-msg global-sim PA 0.887 PA 0.887 PA 0.527 PA 0.823 PA 0.535 learning DW_L2_LR 0.648 DW_L1_RF 0.669 DW_L1_RF 0.178 DW_L2_LR 0.640 DW_L1_RF 0.235 email global-sim Katz 0.611 Norm-CT 0.775 Katz 0.151 Katz 0.583 Katz 0.250 learning GS_Attentional_Avg_LR 0.675 GS_MaxPool_L1_LR 0.722 GS_LSTM_L2_RF 0.252 N2V_Hada_RF 0.666 GS_Mean_L1_RF 0.333 math-flow global-sim PA 0.886 PA 0.869 PA 0.489 PA 0.952 PA 0.476 learning GS_MeanPool_Hada_LR 0.719 GS_MeanPool_Hada_LR 0.714 GS_MeanPool_L1_LR 0.224 GS_MaxPool_L2_LR 0.928 GS_MaxPool_L1_LR 0.291 ast-phys global-sim PA 0.991 PA 0.996 Katz 0.906 PA 0.971 PA 0.857 learning N2V_L2_RF 0.924 GS_MaxPool_L1_RF 0.787 N2V_L2_RF 0.584 N2V_L1_LR 0.971 N2V_L2_RF 0.571 forum global-sim Katz 0.742 Katz 0.779 Katz 0.250 Katz 0.682 Katz 0.346 learning DW_L2_LR 0.625 GS_MaxPool_Hada_LR 0.601 DW_L2_LR 0.127 DW_L2_LR 0.567 DW_L1_RF 0.173 topo global-sim PA 0.879 PA 0.873 CT 0.492 PA 0.779 Norm-CT 0.485 learning GS_MeanPool_Hada_RF 0.448 GS_Mean_Hada_RF 0.880 GS_MeanPool_Hada_RF 0.549 GS_MeanPool_Hada_RF 0.805 GS_Mean_Hada_RF 0.552 pow-grid global-sim Norm-CT 0.616 CT 0.594 CT 0.119 Norm-CT 0.602 -Norm-CT 0.136 learning GS_Mean_L2_RF 0.596 GS_Mean_L2_RF 0.723 GS_MaxPool_L1_LR 0.126 GS_Mean_L1_RF 0.630 N2V_L1_RF 0.150 routers global-sim Norm-CT 0.748 Norm-CT 0.809 Norm-CT 0.288 CT 0.716 CT 0.333 learning GS_MaxPool_Hada_LR 0.720 DW_Hada_LR 0.711 GS_MaxPool_L1_LR 0.216 GS_LSTM_L2_LR 0.641 GS_MaxPool_L1_LR 0.320 boi-yeast global-sim PA 0.762 PA 0.715 CT 0.244 PA 0.933 Norm-CT 0.311 learning DW_L2_LR 0.721 DW_L1_LR 0.696 GS_LSTM_Hada_RF 0.231 DW_L2_LR 0.933 GS_Mean_Avg_RF 0.255 fb global-sim Norm-CT 0.942 Katz 0.944 -NormCT 0.700 PA 0.875 Katz 0.250 learning N2V_L1_LR 0.948 GS_MaxPool_L2_RF 0.935 GS_MaxPool_Hada_RF 0.713 DW_Hada_LR 0.100 DW_Hada_LR 0.625 blog-cat global-sim PA 0.807 PA 0.801 Norm-CT 0.392 PA 0.947 PA 0.391 learning N2V_Avg_RF 0.792 N2V_Avg_LR 0.742 N2V_Avg_RF 0.310 N2V_Avg_RF 0.939 N2V_Avg_RF 0.347
+Appendix C Kendall’s Results for Section 5.1 and Section 5.2
+| NETWORK | AUROC | AUPR | Pr@P | |||
+| balanced | imbalanced | balanced | imbalanced | |||
+| two-hop | CollegeMsg | 0.184 | 0.184 | 0.219 | 0.156 | 0.203 |
+| Email-Eu | 0.272 | 0.319 | 0.472 | 0.294 | 0.526 | |
+| Mathoverflow | 0.339 | 0.386 | 0.414 | 0.326 | 0.488 | |
+| Ast-phys | 0.487 | 0.433 | 0.377 | 0.525 | 0.393 | |
+| lastFm | 0.477 | 0.465 | 0.381 | 0.462 | 0.335 | |
+| IAcontact | 0.027 | 0.080 | 0.038 | 0.039 | 0.075 | |
+| FB-forum | 0.321 | 0.347 | 0.438 | 0.311 | 0.271 | |
+| Topology | 0.785 | 0.783 | 0.771 | 0.723 | 0.742 | |
+| Mooc | 0.744 | 0.743 | 0.698 | 0.678 | 0.741 | |
+| three-hop | CollegeMsg | 0.009 | 0.118 | 0.158 | 0.022 | 0.230 |
+| Email-Eu | -0.030 | 0.085 | 0.167 | -0.144 | 0.181 | |
+| Mathoverflow | -0.020 | -0.071 | 0.018 | 0.139 | 0.060 | |
+| Ast-phys | 0.256 | 0.464 | 0.578 | 0.422 | 0.628 | |
+| lastFm | — | — | — | — | — | |
+| IAcontact | — | — | — | — | — | |
+| FB-forum | 0.189 | 0.208 | 0.283 | 0.203 | 0.334 | |
+| Topology | 0.749 | 0.747 | 0.738 | 0.649 | 0.773 | |
+| Mooc | — | — | — | — | — | |
+| DATASET | AUROC | AUPR | Pr@P | |||
+| balanced | imbalanced | balanced | imbalanced | |||
+| missing LP | clg-msg | 0.275 | 0.203 | 0.241 | 0.316 | 0.207 |
+| -0.059 | -0.012 | -0.155 | -0.091 | 0.007 | ||
+| math-flow | 0.399 | 0.385 | 0.227 | 0.311 | 0.196 | |
+| ast-phys | 0.278 | 0.391 | 0.572 | 0.373 | 0.542 | |
+| lst-fm | — | — | — | — | — | |
+| forum | 0.184 | 0.227 | 0.236 | 0.221 | 0.221 | |
+| topo | 0.233 | 0.316 | 0.302 | 0.176 | 0.241 | |
+| act-mooc | — | — | — | — | — | |
+| pow-grid | -0.144 | 0.084 | 0.173 | -0.108 | 0.189 | |
+| routers | 0.242 | 0.005 | 0.079 | 0.285 | 0.068 | |
+| bio-yeast | 0.216 | 0.260 | 0.362 | 0.316 | 0.302 | |
+| fb | -0.178 | 0.450 | 0.211 | -0.159 | 0.025 | |
+| blog-cat | 0.213 | 0.031 | 0.049 | 0.319 | 0.033 | |
+| future LP | clg-msg-t | 0.580 | 0.589 | 0.548 | 0.578 | 0.489 |
+| email-t | 0.603 | 0.443 | 0.390 | 0.609 | 0.337 | |
+| math-flow-t | 0.387 | 0.437 | 0.467 | 0.391 | 0.445 | |
+| ast-phys-t | 0.660 | 0.705 | 0.677 | 0.508 | 0.653 | |
+| lst-fm-t | 0.112 | -0.021 | -0.095 | 0.222 | -0.131 | |
+| forum-t | 0.476 | 0.467 | 0.511 | 0.488 | 0.406 | |
+| topo-t | 0.363 | 0.391 | 0.360 | 0.326 | 0.361 | |
+| act-mooc-t | 0.439 | 0.589 | 0.573 | 0.413 | 0.594 | |